### v2

# SlideAgent: Hierarchical Agentic Framework for Multi-Page Visual Document Understanding

Abstract
Multi-page visual documents such as manuals, brochures, presentations, and posters convey key information through layout, colors, icons, and cross-slide references. While large language models (LLMs) offer opportunities in document understanding, current systems struggle with complex, multi-page visual documents, particularly in fine-grained reasoning over elements and pages. We introduce SlideAgent, a versatile agentic framework for understanding multi-modal, multi-page, and multi-layout documents, especially slide decks. SlideAgent employs specialized agents and decomposes reasoning into three specialized levels–global, page, and element–to construct a structured, query-agnostic representation that captures both overarching themes and detailed visual or textual cues. During inference, SlideAgent selectively activates specialized agents for multi-level reasoning and integrates their outputs into coherent, context-aware answers. Extensive experiments show that SlideAgent achieves significant improvement over both proprietary (+7.9) and open-source models (+9.8).
SlideAgent: Hierarchical Agentic Framework for Multi-Page Visual Document Understanding
Yiqiao Jin1thanks: Work done as an intern at J.P. Morgan AI Research., Rachneet Kaur2, Zhen Zeng2, Sumitra Ganesh2, and Srijan Kumar1 1Georgia Institute of Technology, 2 J.P. Morgan AI Research {yjin328,srijan}@gatech.edu, zhen.zeng@jpmchase.com, {rachneet.kaur,sumitra.ganesh}@jpmorgan.com https://SlideAgent.github.io/
1 Introduction
Visual documents–from earnings reports and academic lectures to business strategy presentations–are ubiquitous, conveying ideas not only from text, but also from the intricate interplay of layout, icons, visual hierarchy, and cross-page relationships. These documents are central to high-stakes domains such as finance, science, and technology. Accurately interpreting them thus remains a pressing challenge.
Figure 1: When given the full page, the LLM miscounts the number of product mix categories. After isolating the chart, it correctly identifies all eight categories, highlighting the importance of accurate element parsing.
Challenges.
Accurately interpreting these multi-page, multi-modal artifacts remains challenging. Recent advances in multimodal large language models (MLLMs) have accelerated document understanding (Verma et al., 2024), yet three gaps remain: 1) Scalable Fine-Grained Reasoning. State-of-the-art MLLMs process a limited number of images at once (Liu et al., 2023) and tend to treat each page holistically, missing fine-grained, element-level cues required for user queries (Faysse et al., 2024; Tanaka et al., 2025). In Figure 1, an LLM can miscount chart segments on a cluttered page of a slide deck, but correctly identifies all segments once the relevant chart is cropped–highlighting the importance of element-level parsing for latent reasoning abilities. 2) Domain-Specific Visual Semantics. Most MLLMs are pre-trained on natural images (Wu et al., 2024), lacking exposure to domain-specific diagrams, financial charts, or scientific plots. Consequently, they struggle with the specialized language of visual documents (Cho et al., 2024). For example, logos appear on every page, reinforcing brand identities but do not offer additional content. Color schemes encode categorical information (red for losses and green for gains in financial reports). Icons convey abstract concepts (lightbulbs for innovation, arrows for causal relationships); and spatial positioning signals importance (centered elements typically matter more than corner annotations). falter in spatial reasoning Wang et al. (2024b, a), failing to locate visual elements Sharma and Vats (2025); Bhattacharyya et al. (2025); Polak and Morgan (2025). Low-resolution visual encoders in MLLMs (Liu et al., 2023) further miss details such as footnotes or superscripts. 3) Metadata-Free Integration. Many systems  Singer-Vine (2025); huridocs (2025); Rausch et al. (2021) rely on clean metadata–figure locations, hierarchy tags, embedded text layers–that can be unavailable or corrupted in real-world PDF. Users may take screenshots of PDF documents, scan copies of physical documents, export slides or documents as flattened PDFs, upload or share PDFs generated from software that strips out or does not preserve document structure. Recent metadata-free methods (Yu et al., 2025; Tanaka et al., 2025) address these by parsing only visual images without relying on the metadata, despite a performance gap.
This Work.
We present SlideAgent, an LLM-based agentic framework for fine-grained understanding of multi-page, varying-size visual documents. Inspired by the human information processing model (Lang, 2000; Naysmith et al., 2021), SlideAgent employs a hierarchical architecture with specialized agents at three levels: global (document-wide topics), page (page-specific features and cross-page relations), and element (fine-grained components such as charts, figures, and text blocks). During knowledge construction stage, SlideAgent parses layout and generates query-agnostic knowledge at each level. At inference stage, SlideAgent retrieves query-specific knowledge, enabling scalable fine-grained reasoning over relevant pages and elements. We benchmark SlideAgent and baseline models, demonstrating significant performance on both open-source and proprietary models. SlideAgent surpasses its LLM counterpart by 7.9 (proprietary) and 9.8 (open-source), respectively. We also show that SlideAgent enhances spatial reasoning, visual counting, and cross-element understanding, with results that are highly interpretable.
2 Method
Problem Formulation
Given a multi-page visual document $\mathcal{P}=\{p_{1},\ldots,p_{|\mathcal{P}|}\}$ with $|\mathcal{P}|$ pages and a query $q$, the goal is to generate a natural language answer $a=f(q,\mathcal{P})$ by reasoning over relevant visual and textual elements.
Overview
As shown in Figure 2, SlideAgent operates in two stages: 1) Knowledge Construction: Build a hierarchical, query-agnostic knowledge base $\mathcal{K}=\{\mathcal{K}_{g},\mathcal{K}_{p},\mathcal{K}_{e}\}$ capturing global, page, and element knowledge; 2) Retrieval and Question-Answering: Using multi-level retrieval to retrieve query-specific content from $\mathcal{K}$ and synthesize the answer $a$, ensuring both broad contextual understanding and fine-grained reasoning.
Figure 2: SlideAgent generates knowledge about input slide decks in a hierarchical manner at 3 levels: global, page, and element. At each level, specialized agents generate query-agnostic knowledge during knowledge construction, then retrieve and reason over query-specific textual & visual knowledge during inference stage. Sample knowledge $\mathcal{K}$ generated by SlideAgent is in Appendix Figure 7,8,9 and answers generated by the agents are in Figure 5.
2.1 Knowledge Construction Stage
Given a multi-page document, SlideAgent constructs hierarchical knowledge at three levels using specialized agents in a top-down manner.
Global Agent
The global agent $\mathcal{M}_{g}$ generates document-level knowledge $\mathcal{K}_{g}=\mathcal{M}_{g}(\mathcal{P})$, capturing the overall summary, objectives, and narrative flow of the document. This layer establishes overarching themes to support high-level reasoning about the document’s purpose. Since visual documents are often large and LLMs have a limited capacity for processing visuals, SlideAgent samples the first three pages to generate $\mathcal{K}_{g}$ (example in Appendix Figure 7).
Page Agent
For each page $p_{i}\in\mathcal{P}$, the page agent generates page-level knowledge $\mathcal{K}_{p}$ in a sequential manner, conditioned on the page’s visual content $v_{i}$, the global knowledge $\mathcal{K}_{g}$, and the knowledge from the preceding page $\mathcal{K}_{p}^{i-1}$:
$\mathcal{K}_{p}^{i}=\mathcal{M}_{p}(v_{i},\mathcal{K}_{g},\mathcal{K}_{p}^{i-1}),i\in[1,|\mathcal{P}|].\quad\mathcal{K}_{p}^{0}=\emptyset$ (1)
The complete page-level knowledge $\mathcal{K}_{p}=\bigcup_{i=1}^{|\mathcal{P}|}\mathcal{K}_{p}^{i}$ provides an intermediate representation that captures page-specific content while linking them to the global context (sample in Figure 8). To ensure comprehensive understanding of all pages, $\mathcal{K}_{p}$ is subsequently used to refine $\mathcal{K}_{g}$.
Element Agent
LLMs often struggle with spatial reasoning over visual documents Wang et al. (2024b, a), failing to locate visual elements Sharma and Vats (2025); Bhattacharyya et al. (2025); Polak and Morgan (2025). To address this, our element agent integrates external tools to explicitly capture the spatial and structural information of each page.
At the finest granularity, the element agent decomposes each page $p_{i}$ into a set of elements using a layout parsing pipeline $f:v_{i}\rightarrow\{(i,e_{j},b_{j},t_{j})\}_{j=1}^{M_{i}}$, where each element is represented by its page index $i$, verbatim text $e_{j}$, bounding box coordinates $b_{j}$, and element type $t_{j}$. $f$ integrates text detection, layout detection, and element classification, followed by post-processing to merge fragmented elements (Appendix A.1).
For each detected element $e_{j}$, the agent consumes the annotated visuals and metadata to generate element-level knowledge:
$\mathcal{K}_{e}^{j}=\mathcal{M}_{e}(i,v_{i},e_{j},b_{j},t_{j},\mathcal{K}_{g},\mathcal{K}_{p}^{i}).$ (2)
$\mathcal{K}_{e}^{j}$ consists of the element’s semantic role, functional purpose, and its relation to the slide page (example in Figure 9). This design allows SlideAgent to reason consistently across diverse visual content while preserving spatial relationships for document understanding.
2.2 Inference Stage
Query Classification
Different queries require different perspectives to answer effectively. For instance, queries about global understanding, like “What is the overall theme of this presentation?” requires a broad overview from a macroscopic perspective and activates only the global agent. In contrast, a fact-based query, like “What is the revenue on slide 3?” requires detailed, slide-specific information and requires both the page and element agents. Leveraging too many agents may increase computation or introduce noise. Thus, the agent orchestrator first attempts to classify each query $q$ into one of 4 predefined categories, such as global understanding, fact-based direct queries, multi-hop reasoning, and layout & visual relationships. Each type corresponds to a question-answering strategy and a targeted set of agents (Appendix Table 7). For instance, global understanding queries activate only the global agent, as they require a broad overview, while fact-based queries trigger both the page and element agents to retrieve detailed, slide-specific information. If none of the predefined categories apply, the agent defaults to the “unknown” category, which activates all agents.
Subquery Generation and retrieval
The original query $q$ is usually short and can lead to noisy retrieval. Using $q$, SlideAgent generates subqueries $\hat{Q}$ targeting key entities in the query. For example, for the query ‘Which country has the smallest GWP in the time range described by the donut chart,’ the model generates subqueries related to keywords such as country, donut chart, and GWP. $q$ and $\hat{Q}$ are concatenated to jointly retrieve the top-$k_{\ell}$ pages $\hat{\mathcal{P}}$ and the top elements $\hat{\mathcal{E}}$ along with their page / element knowledge $\mathcal{K}_{p}$ and $\mathcal{K}_{e}$. Note that the retriever can include simple sparse retrievers such as BM25 Robertson et al. (2004), dense retriever such as SFR Meng et al. (2024), GTE (Li et al., 2023) and Linq (Kim et al., 2024), and multimodal retriever such as COLPALI Faysse et al. (2024) and VisRAG (Yu et al., 2025).
Answer Generation and Synthesis
The system guides structured reasoning through hierarchical context understanding to generate $h_{g},h_{p},h_{e}$, which contain both the agent’s answer and its reasoning. The global context is processed to generate $h_{g}=f_{g}(\mathcal{K}_{g},\mathcal{P}_{s},q)$, which captures the overall document-level context. The page-level agent then derives from the retrieved pages and corresponding knowledge: $h_{p}=f_{p}(\mathcal{K}_{p},\mathcal{R}_{p}(\mathcal{P},{\{q\}\cup\hat{Q}}),h_{g})$. At the element-level, the visual input is annotated with bounding boxes $\{b_{i}\}$, which are then processed by the element agent to generate $)h_{e}=f_{e}(\mathcal{R}^{\prime}_{e}(\mathcal{E},\{q\}\cup\hat{Q}),\mathrm{Annot}(\hat{V}$)), capturing the detailed visual and textual cues.
If all agents agree according to answer matching (Appendix B.2), or only one agent is activated, the answer is taken directly from the activated agent. Otherwise, the answer synthesizer $\phi(\cdot)$ combines the reasoning from all agents and visuals from the retrieved pages to generate the final answer:
$).a=\phi(h_{g},h_{p},h_{e},\{v_{i}:p_{i}\in\mathcal{R}(\hat{\mathcal{P}}$)\}). (3)
3 Experiments
Model SlideVQA TechSlides FinSlides
Overall Num F1 Overall Num F1 Overall Num F1
Multimodal LLMs (Type 1)
Gemini 2.0 75.0 71.3 79.8 50.4 67.6 41.6 70.8 70.6 77.8
Gemini 2.5 83.8 78.3 91.8 51.1 71.4 41.2 76.2 75.8 100.0
Gemini 2.5-lite 71.2 60.8 87.0 47.3 58.1 41.9 57.0 56.6 68.3
Claude 4.1 78.4 74.3 82.3 61.0 81.4 52.3 56.5 54.8 73.3
Claude 3.5 62.5 68.3 54.6 52.5 80.2 39.5 48.5 49.5 29.6
GPT-4o 77.0 72.1 84.0 63.4 78.3 53.9 80.0 80.8 62.1
Multimodal RAG (Type 2) and Agentic Methods (Type 3)
COLPALI 78.8 73.7 83.4 64.1 73.2 54.5 80.9 81.5 62.7
VisRAG 78.2 73.1 85.4 64.7 72.6 54.7 79.2 81.1 75.8
VDocRAG 80.0 75.0 87.8 67.0 80.5 57.0 83.5 83.8 64.2
ViDoRAG 81.1 76.4 88.1 68.7 78.2 59.4 82.2 83.3 65.1
SlideAgent 84.9 80.4 90.5 70.9 82.5 66.2 85.5 85.9 79.6
Impr. +7.9 +8.3 +6.5 +7.5 +4.2 +12.3 +5.5 +5.0 +17.5
Table 1: Performance comparison of proprietary models on SlideVQA, TechSlides, and FinSlides. All baseline methods use GPT-4o for answer generation. Improvements over the base model GPT-4o is shown in green. The best and second best performance are highlighted in bold and underlined. SlideAgent outperforms all baseline methods (Type 2/3) sharing the same base model (GPT-4o), and even outperforms stronger raw LLMs (e.g. Gemini-2.5).
Datasets
We evaluate SlideAgent and baselines on two tasks: 1) multi-page understanding, using datasets such as SlideVQA (Tanaka et al., 2023), TechSlides, and FinSlides (Wasserman et al., 2025); and 2) single-page understanding, using InfoVQA (Mathew et al., 2022). Details of the datasets are in Appendix C.1.
Models.
We benchmark SlideAgent against 3 types of baselines: 1) Multimodal LLMs: 15 LLMs from 8 model families, including proprietary models (GPT-4o (OpenAI, 2025), Gemini (Anil et al., 2023), Claude (Anthropic, 2025)) and open-source models (Llama-3.2 (Grattafiori et al., 2024), InternVL3 (Chen et al., 2024), Phi-3 (Abdin et al., 2024), Qwen2.5-VL (Bai et al., 2025)); 2) Multimodal RAG Methods: VisRAG Yu et al. (2025), VDocRAG Tanaka et al. (2025), and COLPALI Faysse et al. (2024); 3) Multi-agent Systems: ViDoRAG Wang et al. (2025a). We evaluate SlideAgent using two backbone LLMs, including both proprietary models (GPT-4o (OpenAI, 2025)) and open-source models (InternVL3-8B (Chen et al., 2024)). These models are chosen for their widespread use in state-of-the-art QA systems Yu et al. (2025); Jin et al. (2025); Cho et al. (2024). The parameter sizes, knowledge cutoff dates, and release dates are in Appendix Table 8. We use the text-based retriever SFR (Meng et al., 2024) due to its strong efficiency-performance Cheng et al. (2024); Jin et al. (2025). For models restricted to single-image input (e.g. LLaVA-v1.5 (Liu et al., 2023)), we concatenate the top 3 retrieved images into one following previous work Yu et al. (2025). In settings where ground-truth pages are available, we provide them as input to all models. Otherwise, each model receives as many retrieved pages up to its input capacity.
Figure 3: Performance comparison among variants of SlideAgent with base model GPT-4o.
Metrics.
We evaluate the performance of SlideAgent in end-to-end question answering. For questions asking about numeric values, we extract, standardize, and compare the prediction and the ground-truth in various formats, including percentages, decimals, integers, and word-based representations (e.g., “three”, “thousand”, “million”). Numbers are normalized to a unified format (e.g., ‘17k’ $\rightarrow$ ‘17000’, ‘2.5 million’ $\rightarrow$ ‘2500000’, ‘97%’ $\rightarrow$ ‘0.97). Otherwise, we use F1-score to evaluate the lexical overlap between predicted and ground-truth answers. Both answers are normalized, tokenized, and preprocessed by removing stopwords and punctuation before metric calculation. For ranking, we use MRR, Hit@k, and nDCG@k (Appendix C.2)
Settings
We evaluate SlideAgent under two realistic settings: 1) End-to-End Performance. The model must first retrieve relevant pages before answering the query, reflecting real workflows where users query long visual documents –analysts navigating 100-page earnings presentation, students reviewing lecture slides, or engineers inspecting multi-page technical specifications. This setting measures the end-to-end capability in retrieval, spatial reasoning, and layout understanding. 2) Performance with Ground-truth Pages. The model is directly given the page(s) containing the answer, isolating reasoning from retrieval. This mirrors cases where context is known–e.g., a product manager reviewing a linked design slide–and primarily focuses on reasoning and layout comprehension.
Model Overall Num F1
Raw LLMs
Gemini 2.0 72.3 66.1 81.3
Gemini 2.5 86.9 81.1 95.4
Gemini 2.5 lite 71.7 62.6 87.0
Claude 4.1 36.3 35.4 38.9
Claude 3.5 31.5 30.8 37.0
GPT-4o 69.0 59.3 90.5
Multimodal RAG and Agentic Methods
ViDoRAG 71.2 60.5 90.7
SlideAgent 79.6 69.9 94.1
Impr. +10.6 +10.5 +3.6
Table 2: Performance comparison among Gemini, Claude, GPT-4o, ViDoRAG, and SlideAgent models.
3.1 End-to-End Performance
Tables 1/2/3 and Appendix Table 5 demonstrate the end-to-end performance of SlideAgent across proprietary and open-source models.
Model SlideVQA TechSlides FinSlides
Overall Num F1 Overall Num F1 Overall Num F1
Multimodal LLMs (Type 1)
Llama 3.2 11B 42.9 43.3 42.3 41.4 52.5 36.2 23.3 23.2 26.2
Phi3 72.3 61.8 90.6 59.4 60.0 59.1 48.8 48.5 64.3
Qwen2.5 7B 79.5 70.5 94.3 59.3 52.5 65.9 53.6 52.5 85.7
Qwen2.5 32B 79.2 71.1 92.2 67.5 87.5 60.6 57.4 56.6 87.5
LLaVA 1.5 7B 36.8 22.4 79.0 23.3 12.5 38.7 10.7 10.1 16.6
LLaVA 1.5 13B 44.9 25.1 81.8 28.1 17.5 45.0 20.6 16.5 36.7
LLaVA 1.6 7B 50.9 37.3 82.6 34.4 37.5 32.2 12.2 12.1 17.8
LLaVA 1.6 13B 16.7 10.2 81.5 45.2 40.0 49.1 32.0 31.3 64.3
InternVL3 8B 63.0 56.5 74.1 55.4 57.5 54.4 49.8 49.5 64.3
Multimodal RAG (Type 2) and Agentic Method (Type 3)
COLPALI 63.4 56.7 73.8 57.1 60.9 55.2 50.4 49.3 65.7
VisRAG 63.6 56.5 75.5 56.8 57.7 55.4 51.1 49.6 65.2
VDocRAG 65.2 59.7 77.0 59.2 60.7 58.3 51.8 50.1 65.9
ViDoRAG 68.8 61.9 77.3 61.4 61.9 59.3 52.7 55.4 66.6
SlideAgent 72.7 68.2 79.4 63.1 78.0 61.7 63.3 62.8 68.3
Impr. +9.8 +11.7 +5.4 +7.7 +20.5 +2.3 +13.5 +13.3 +4.0
Table 3: Performance comparison of various models on SlideVQA, TechSlides, and FinSlides datasets. SlideAgent outperforms all baseline methods (Type 2/3) sharing the same base model (GPT-4o).
Consistent Improvements across Architectures
For both proprietary and open-source models, SlideAgent consistently outperforms all baseline methods (Type 2/3) and the base models across all metrics. For proprietary models (Table 1), SlideAgent improves overall accuracy by +7.9, numeric reasoning by +8.3, and lexical overlap by +6.5 on SlideVQA. Similar trends are observed for TechSlides (+7.5 overall; +12.3 F1) and FinSlides (+5.5 overall; +17.5 F1), indicating robust performance across diverse domains. For open-source models, the advantage remains pronounced: SlideAgent outperforms InternVL3-8B by +9.8 overall (+11.7 numeric), showing that our hierarchical, multi-agent design generalizes well across LLMs.
Comparison with Multimodal LLMs
SlideAgent consistently achieves the best or second-best performance among proprietary models (Tables 1 and  11). Notably, although the base model slightly lags behind stronger LLMs such as Gemini 2.5 in raw capability (e.g. 83.8 for Gemini 2.5 vs. 77.0 for GPT-4o in Table 1), SlideAgent’s structured reasoning pipeline fills this gap (84.9 overall for SlideAgent). For open-source models, SlideAgent achieves better performance across all models except for the Qwen2.5 family. While strong base models such as Gemini-2.5 and Qwen2.5 already exhibit advanced multimodal comprehension, SlideAgent is model-agnostic in nature and can be directly applied to these models to further enhance performance.
Figure 4: Accuracy of SlideAgent and base models (GPT-4o / InternVL3-8B) on different query types.
Performance across Query Types
Figure 4 presents the performance breakdown across query types, as defined in Table 7. SlideAgent improves question-answering across diverse query types, especially in multi-hop reasoning and visual/layout questions. The greatest improvement occurs on case 3 (multi-hop reasoning), with a 9.8 point improvement (67.4 to 77.2). This means explicitly guiding the model’s reasoning using generated knowledge $(\mathcal{K}_{g},\mathcal{K}_{p},\mathcal{K}_{e})$ significantly improves reasoning capabilities. A notable 7.7-point improvement (66.7 $\rightarrow$ 74.4) is also observed in visual/layout reasoning, demonstrating the benefit of fine-grained element-level reasoning and retrieval that multimodal LLMs hardly achieve. Both SlideAgent and its counterpart perform well on Case 2 (Fact-based Direct Queries), with only a modest 2.1-point improvement, reflecting the model’s ability to handle page-level reasoning with little space for further enhancement.
3.2 Performance with Ground-truth Pages
When ground-truth pages are provided (Table 11/12), the performance gap narrows. As all models receive the exact pages with the answers, noise introduced in retrieval is eliminated. Under this oracle setting, SlideAgent still improves over the base model (+7.7 overall and +12.5 numeric on SlideVQA), demonstrating the effectiveness of element-level retrieval.
Text-based Retrievers MRR Recall@1 nDCG@1 Recall@3 Hit@3 nDCG@3
BM25 Robertson et al. (2004) 59.0 51.5 52.0 63.0 65.3 57.7
w/ SA 63.9 +4.9 54.9 +3.4 56.6 +4.6 67.1 +4.1 68.8 +3.5 62.6 +4.9
BGE Xiao et al. (2023) 70.1 56.1 60.9 75.8 78.1 69.2
w/ SA 72.3 +2.2 58.4 +2.3 63.2 +2.3 77.4 +1.6 80.3 +2.2 71.3 +2.1
SFR Meng et al. (2024) 70.1 56.1 60.9 75.8 78.1 69.2
w/ SA 76.5 +6.4 59.9 +3.8 69.4 +8.5 77.3 +1.5 81.8 +3.7 73.2 +4.0
Multimodal Retrievers MRR Recall@1 NDCG@1 Recall@3 Hit@3 NDCG@3
SigLIP2 Tschannen et al. (2025) 26.7 15.9 18.0 31.0 34.0 25.0
w/ SA 28.0 +1.3 16.3 +0.4 18.0 +0.0 32.3 +1.3 35.5 +1.5 26.1 +1.1
COLPALI Faysse et al. (2024) 82.1 68.2 75.5 78.9 84.0 77.4
w/ SA 82.9 +0.8 70.4 +2.2 76.2 +0.7 88.6 +9.7 90.1 +6.1 83.1 +5.7
VisRAG Yu et al. (2025) 76.0 63.3 68.6 82.3 84.1 76.0
w/ SA 79.7 +3.7 66.3 +3.0 71.6 +3.0 85.5 +3.2 87.7 +3.6 79.4 +3.4
Table 4: Artifacts generated by SlideAgent, particularly $\hat{Q}$ and $\mathcal{M}_{p}$, improves page-level retrieval performance, especially for text-based retrievers.
3.3 Effectiveness of Knowledge Construction
We evaluate whether hierarchical knowledge representations ($\mathcal{K}$) improve retrieval beyond end-to-end QA. Specifically, we test page-level retrieval using generated subqueries $\hat{Q}$ and page knowledge $\mathcal{K}_{p}$ from SlideAgent. Table 4 shows consistent gains across both text-based retrievers (BM25 Robertson et al. (2004), BGE Xiao et al. (2023), SFR Meng et al. (2024)) and multimodal retrievers (COLPALI Faysse et al. (2024), VisRAG Yu et al. (2025), SigLIP2 Tschannen et al. (2025)).
Text-based Retrievers Show Largest Gains.
Structured agent outputs substantially enhance text-based retrievers. SFR achieves the largest gains (+6.4 MRR, +8.5 nDCG@1), showing that page-level knowledge $\mathcal{K}_{p}$ provides richer semantic signals than raw OCR, especially useful in the absence of the vision modality. Even sparse retrievers like BM25 improves (+4.9 MRR), as lexical matching benefits from structured page descriptions, which integrates multimodal cues that pure text extraction often miss. Notably, despite much lower computational costs, text-based retrievers already rival multimodal LLM-based methods, justifying our choice of SFR as the base retriever.
Multimodal Retrievers Exhibit Smaller but Consistent Gains.
COLPALI Faysse et al. (2024) improves slightly in ranking quality (+0.8 MRR) but substantially in coverage (+9.7 Recall@3), indicating that structured subqueries help it surface more relevant pages even if they are not ranked at the very top. VisRAG Yu et al. (2025) achieves a larger ranking boost (+3.7 MRR), suggesting that multimodal LLMs still benefit from textual guidance in aligning queries to page content. SigLIP2 shows minimal gains (+1.3 MRR), likely because it is optimized for natural image domains and transfers less effectively to document-style inputs.
Figure 5: Example answers generated by SlideAgent. Agents at different levels work together to provide comprehensive responses.
3.4 Ablation Studies
We analyze each design choice by removing global agent (w/o G), page agent (w/o P), element agent (w/o E), and subquery generation (w/o S). Figures 3/6 summarize results across proprietary (GPT-4o) and open-source (InternVL3-8B) backbones.
Page-level Reasoning is Critical
Removing the page agent causes the steepest degradation–GPT-4o drops –6.3 overall (–9.5 numeric) and InternVL3-8B –8.8 overall. The page agent integrates global themes $\mathcal{K}_{g}$ and sequential context $\mathcal{K}_{p}^{i-1}$ (Eq. 1), enabling cross-slide coherence and multi-hop reasoning. Its absence breaks this structural bridge between global context and page details. Results for SlideAgent w/o E, w/o G, and w/o S are in Appendix D.1.
3.5 Qualitative Answer Analysis
Figure 5 shows a qualitative example of answer generated by SlideAgent to illustrate how SlideAgent leverages hierarchical knowledge for reasoning.
Given the slide deck, SlideAgent builds hierarchical knowledge $\mathcal{K}=\{\mathcal{K}_{g},\mathcal{K}_{p},\mathcal{K}_{e}\}$ (Figures 79). Because the query lacks global/page cues (e.g. “on page 4”), the agent orchestrator activates all agents. The global agent performs deck-level triage and nominates the pages directly following page 2’s summary as related to “cause/effect region.” The page agent pinpoints Page 4: Wealth Management–The Cause, whose description states that listed problems leads to Business Under-Performance. Element ($\mathcal{K}_{e}$) grounds the answer by parsing the flowchart on Page 4, extracting the node whose verbatim text matches the query and following its directed edge to the target node labeled Business Under-Performance. The answer synthesizer fuses these signals to return Business under-performance with explicit provenance (Page 4, flowchart edge from the queried node). This layering is essential: without global triage, search drifts; without page focus, we may match the phrase but miss causality; without element grounding we rely on summaries and risk speculation. Combining all three yields robust reasoning: the global agent provides thematic scope, the page agent narrows candidates, and the element agent confirms answers with visual grounding. This layered verification produces accurate, confident responses.
4 Related Work
In this section, we discuss relevant works in visual document understanding and Multimodal LLMs. A more detailed description is in Appendix B.
Visual Document Understanding relied on computer vision pipelines combining OCR, layout parsing, and heuristics to extract semantics Shilman et al. (2005); Bhowmik (2023); Gao et al. (2019); Kieninger (1998); Smith (2007). Though effective, these methods were brittle to noisy and visually rich inputs. Recent models such as BERTgrid Denk and Reisswig (2019), LayoutLM Xu et al. (2020b, a), LayoutT5 Tanaka et al. (2021), and TILT Powalski et al. (2021) jointly encode texts, layouts, and visuals through multimodal pretraining–laying the foundation for LLM document understanding. Visual Question Answering based on Multi-page Document requires layout comprehension and long-context reasoning. Earlier studies focused on segmentation Haurilet et al. (2019) and generation Sun et al. (2021); Fu et al. (2022) perspectives, while recent work focus on multi-hop, numerical, or commonsense reasoning capabilities Tanaka et al. (2023); Mathew et al. (2022); Ma et al. (2024). Retrieval pipelines such as ColPali Faysse et al. (2024), VisRAG Yu et al. (2025), and VDocRAG Tanaka et al. (2025) combine retrieval and reasoning for multi-page comprehension, emphasizing the need for fine-grained element-level reasoning.
General-Purpose Multimodal LLMs such as GPT-4/4o Achiam et al. (2023); Hurst et al. (2024), Gemini Anil et al. (2023), LLaVA Liu et al. (2023), and InternVL3 Zhu et al. (2025) have advanced visual reasoning Shao et al. (2024); Yang et al. (2023). Yet, trained largely on natural images Wu et al. (2024), these models struggle with visual documents such as slides, which require precise grounding of heterogeneous elements (charts, tables, icons).
5 Conclusion
We present SlideAgent, a hierarchical agentic framework that leverages specialized agents at multiple levels for fine-grained infographics understanding, particularly for slide decks. SlideAgent demonstrates significant performance gains and strong generalizability across both proprietary and open-source models.
6 Limitations
We did not evaluate element-level retrieval due to the lack of corresponding annotations in existing datasets and the difficulty of defining precise boundaries among elements. Benchmarks in visual document understanding could include such annotations to enable finer-grained analysis.
Retrieval Method.
Our framework supports both textual and multimodal retrievers. We primarily adopt text-based retrieval for efficiency, though future work could explore multimodal retrievers or domain-specific strategies tailored to slide decks.
7 Ethical Considerations
Data Privacy, Consent, and Intellectual Property.
Visual documents such as those processed by SlideAgent may contain sensitive business or personal data. Ensuring compliance with privacy regulations (e.g., GDPR, CCPA) and obtaining appropriate consent are essential. Organizations should also respect intellectual property rights and establish policies that balance knowledge sharing with fair use.
Content Reliability and User Responsibility.
Like other LLM-based systems, SlideAgent may inherit biases or produce inaccurate content. While it enhances question answering, outputs should be verified by human judgment, particularly in sensitive or high-stakes scenarios. Clear user guidance and validation practices can help ensure responsible use.
Disclaimer
This paper was prepared for informational purposes by the Artificial Intelligence Research group of JPMorgan Chase & Co and its affiliates ("J.P. Morgan") and is not a product of the Research Department of J.P. Morgan. J.P. Morgan makes no representation and warranty whatsoever and disclaims all liability, for the completeness, accuracy or reliability of the information contained herein. This document is not intended as investment research or investment advice, or a recommendation, offer or solicitation for the purchase or sale of any security, financial instrument, financial product or service, or to be used in any way for evaluating the merits of participating in any transaction, and shall not constitute a solicitation under any jurisdiction or to any person, if such solicitation under such jurisdiction or to such person would be unlawful.
Appendix A Details Method
A.1 Element Merging
Raw output, especially text spans from the element processing pipeline (Section 2.1) can fragment coherent texts into multiple parts, creating challenges for downstream understanding. We thus adopt a graph-based merging algorithm to reconstruct semantically coherent text blocks while preserving spatial layout.
Distance-Based Adjacency
Our merging criterion is based on minimum distance between bounding boxes. For two boxes $b_{i}=(x_{1}^{i},y_{1}^{i},x_{2}^{i},y_{2}^{i})$ and $b_{j}=(x_{1}^{j},y_{1}^{j},x_{2}^{j},y_{2}^{j})$, the minimum distance is computed from the horizontal and vertical distances $d_{h}$ and $d_{v}$:
$d_{\text{min}}(b_{i},b_{j})=\sqrt{d_{h}^{2}+d_{v}^{2}}$
where $d_{h}=\min(|x_{2}^{i}-x_{1}^{j}|,|x_{2}^{j}-x_{1}^{i}|)$ if boxes don’t overlap horizontally, and $d_{h}=0$ otherwise. Similarly, $d_{v}=\min(|y_{2}^{i}-y_{1}^{j}|,|y_{2}^{j}-y_{1}^{i}|)$ if boxes don’t overlap vertically, and $d_{v}=0$ otherwise.
Two boxes are considered adjacent if $d_{\text{min}}(b_{i},b_{j})\leq\tau$ where $\tau$ is a threshold we choose to be 15 pixels. The results are fed into the following graph-based component detection.
Notation Description
$p_{i},v_{i}$ A slide page and its visual content
$\mathcal{P}=\{p_{1},p_{2},\dots\}$ Set of pages in the slide deck
$q$ User query
$\hat{Q}$ Set of subqueries generated from the original query $q$
$a$ The final answer generated by SlideAgent
$e_{j}$ Individual element in a slide, consisting of visual content, bounding box, and type
$b_{j}$ Bounding box coordinates of an element within a slide
$t_{j}$ Type of the element (e.g., text, chart, image, table)
$\mathcal{K}_{g},\mathcal{K}_{p},\mathcal{K}_{e}$ Global, page, and element knowledge
$\mathcal{M}_{g},\mathcal{M}_{p},\mathcal{M}_{e}$ Global, page, and element agent
$\mathcal{R}$ Retrieval function that fetches relevant pages or elements based on subqueries
$\phi(\cdot)$ Answer synthesizer that combines reasoning from all agents to generate the final answer $a$
[…]
$h_{g},h_{p},h_{e}$ Answers and reasoning from global, page, and element agents
Table 6: Mathematical notation used throughout the paper.
Case Characteristics Example Agents
1-Global Understanding Asks about the overall theme, purpose, or general summary of the entire presentation. "What is the main topic of the presentation?", "What is this deck about?" Global
2-Fact-based Direct Query Asks for specific facts, data, or information from particular slides. "What is the revenue reported on slide 7?", "Which slide shows the product roadmap?" Page, Element
3-Multi-hop Reasoning Requires comparing information across multiple slides or elements. "Compare revenues in slide 5 and slide 10", "How do Sweden and Denmark compare?" Global, Page, Element
4-Layout / Visual Relationship Asks about visual relationships, positioning, or layout elements. "What does the diagram below the table on slide 12 illustrate?", "Is the color red used to denote negative performance?" Element
5-Uncertain If the query is unclear, use all agents to answer the query. \ Global, Page, Element
Table 7: SlideAgent’s query classification that determines which hierarchical agents to activate.
Appendix B Extended Related Work
B.1 Document and Infographics Understanding
Early research in document and infographic understanding predates multimodal LLMs and was dominated by computer vision pipelines. Classical approaches typically combined OCR, layout parsing, and heuristic rules to extract semantics from text-rich visual content such as forms, tables, and scientific figures Shilman et al. (2005); Bhowmik (2023); Gao et al. (2019); Kieninger (1998); Smith (2007). While these methods enabled structured information extraction, they were brittle to domain shifts and noisy scans.
With the rise of deep learning, several specialized benchmarks spurred progress in document visual question answering (VQA). Datasets such as DocVQA Mathew et al. (2021), DocCVQA Tito et al. (2021), FUNSD Jaume et al. (2019), and DocBank Li et al. (2020) emphasized reasoning over diverse document layouts, scanned forms, and large-scale document collections. ChartQA Masry et al. (2022) further extended this line by introducing reasoning over structured figures like bar and line charts. On the modeling side, early work on text-centric VQA demonstrated that reading and interpreting embedded text was crucial Singh et al. (2019); Mishra et al. (2019). To better represent visually rich documents, approaches such as BERTgrid Denk and Reisswig (2019) and LayoutLM Xu et al. (2020b) proposed contextualized embeddings that jointly encode textual content and 2D spatial layout. This line was extended by LayoutLMv2 Xu et al. (2020a), LayoutT5 Tanaka et al. (2021), and TILT Powalski et al. (2021), which integrated stronger visual features and multimodal pretraining objectives. These advances have achieved impressive results on single-image document VQA tasks by unifying textual, layout, and visual signals, laying the foundation for more recent LLM-based systems.
B.1.1 Slide Deck Understanding
Building on document and infographic understanding, research has also explored presentation slides, which pose unique challenges due to their combination of dense text, figures, and multi-page structure. Early efforts addressed component-level analysis, such as object segmentation on slide pages Haurilet et al. (2019), or generation tasks like creating slides from research papers Sun et al. (2021); Fu et al. (2022). More recently, benchmarks such as MMLongBench Ma et al. (2024) have emphasized long-context processing over lengthy multi-page PDFs, highlighting the scalability issues that arise when moving from single documents to multi-slide collections.
B.1.2 Slide Visual Question Answering
Slide visual question answering (Slide VQA) emerges as a prominent direction within this broader area, aiming to automatically interpret presentation slides to answer natural-language queries. Early work such as SlideVQA Tanaka et al. (2023) introduced multi-modal slide decks paired with questions requiring single-hop, multi-hop, and numerical reasoning across slides. This line was extended by InfoVQA Mathew et al. (2022), which targeted information-centric slides and documents, often involving arithmetic and commonsense reasoning. More recent benchmarks such as TechSlides and FinSlides Wasserman et al. (2025) from the REAL-MM-RAG suite emphasize domain-specific contexts—technical and financial presentations that integrate text, figures, and tables.
At the same time, multi-image and multi-context evaluation benchmarks such as MIBench Liu et al. (2024) and MC-Bench Xu et al. (2024) highlight the unique reasoning challenges in multi-slide and multi-panel scenarios. Retrieval-based methods such as ColPali Faysse et al. (2024) improve slide-grounded search and QA, while pipeline approaches like VisRAG Yu et al. (2025) and VDocRAG Tanaka et al. (2025) integrate retrieval with reasoning to support multi-page comprehension. Collectively, these datasets and methods underscore both the promise and complexity of slide understanding, while leaving open the question of whether element-level reasoning—beyond page-level retrieval—can substantially improve performance.
B.1.3 General Purpose LLMs
General-purpose large language models (MLLMs) such as GPT-4 Achiam et al. (2023), GPT-4o Hurst et al. (2024), Gemini Anil et al. (2023), LLaVA Liu et al. (2023), InternVL3 Zhu et al. (2025), and Visual CoT Shao et al. (2024) have significantly advanced visual reasoning across a wide range of tasks. Visual prompting methods like Set-of-Marks Yang et al. (2023) further enhance reasoning by augmenting inputs with structured visual annotations. While these open-source and closed-source MLLMs form the foundation for many comparisons, they often struggle when applied to slides, where precise grounding and fine-grained interpretation are required to parse heterogeneous elements such as charts, tables, and icons.
Yet key challenges remain. Domain-specific visual semantics are often overlooked, as LLMs trained primarily on natural images Wu et al. (2024) struggle to capture the specialized conventions of slides and infographics–for instance, repetitive logos, color-coded encodings, or abstract symbolic icons. Metadata-dependent integration is similarly fragile: real-world slides and PDFs frequently lack clean structural annotations, and scanned or flattened exports strip away layout cues, rendering metadata-reliant systems brittle huridocs (2025); Rausch et al. (2021). Finally, scalable fine-grained reasoning capabilities remains limited Jin et al. (2024). Current models can process only a small number of images at once Liu et al. (2023), while retrieval typically occurs at the page level Faysse et al. (2024); Tanaka et al. (2025), overlooking element-level reasoning.
Our work builds on these developments by introducing a metadata-free, slide-specialized agentic framework. Unlike prior page-level approaches, our method constructs hierarchical representations and employs retrieval-then-reasoning at the global (deck-level), page (slide-level), and element (fine-grained component) layers. This design enables robust comprehension across diverse slide domains and opens the door to systematic evaluation of whether element-level reasoning provides measurable improvements over page-level retrieval alone.
Adjacency Graph Construction. An undirected graph $G=(V,E)$ is constructed, where vertices $V=\{1,2,\ldots,|B|\}$ represent the bounding boxes for the textual elements, and $E$ edges connect adjacent boxes. An edge $(i,j)\in E$ exists if $d_{\text{min}}(b_{i},b_{j})\leq\tau$.
Connected Component Detection. We identify coherent text spans using depth-first search (DFS) to find connected components. Each component $C_{k}=\{i_{1},i_{2},\ldots,i_{m}\}$ represents indices of boxes that should be merged into a single text block.
Spatial Merging and Text Concatenation. For each component $C_{k}$, we compute the unified bounding box:
$B_{k}=\left(\min_{i\in C_{k}}x_{1}^{i},\min_{i\in C_{k}}y_{1}^{i},\max_{i\in C_{k}}x_{2}^{i},\max_{i\in C_{k}}y_{2}^{i}\right)$
The final text is given by concatenating text in the reading order. This is achieved by sorting boxes within each component using lexicographic ordering: primary sort by vertical position ($y_{1}$), secondary sort by horizontal position ($x_{1}$).
This approach effectively reduces element fragmentation while preserving spatial relationships, enabling more accurate element-level descriptions for our hierarchical framework. Algorithm 1 provides the complete implementation details.
B.2 Answer Matching
To determine answer equivalence, i.e. if two agents agree upon their opinions, we do fuzzy string matching between each pair of answers $a_{1},a_{2}$ using Normalized Levenshtein Similarity (NLS):
$\text{NLS}(a_{1},a_{2})=\left(1-\frac{\text{editdistance}(a_{1},a_{2})}{\max(|a_{1}|,|a_{2}|)}\right),$ (4)
where $\text{editdistance}(a_{1},a_{2})$ computes the Levenshtein distance (Lcvenshtcin, 1966) between the answers after tokenization. Two answers are considered equivalent if $\text{NLS}(a_{1},a_{2})\geq 0.75$.
Appendix C Experimental Setup
C.1 Dataset Descriptions
  • SlideVQA (Tanaka et al., 2023) is a multi-modal, multi-image VQA dataset consisting of $\geq$ 2600 slide decks that require single-hop, multi-hop, and numerical reasoning. Each slide deck corresponds to multiple questions. The license is availble at  [^1] https://github.com/nttmdlab-nlp/SlideVQA/blob/main/LICENSE.
  • InfoVQA (Mathew et al., 2022) is a dataset that feature questions that require reasoning and arithmetic skills. It contains over 30,000 questions with over 5,400 images. The dataset is released under CC-BY license.
  • TechSlides and FinSlides (Wasserman et al., 2025) are from the REAL-MM-RAG benchmark, which comprises slides and documents with text, figures, tables, and images, requiring systems to handle combined textual and visual data. The dataset is released under the CDLA-Permissive-2.0 license.
C.2 Metrics
For ranking evaluation, we adopt MRR, Hit@k, and nDCG@k. Mean Reciprocal Rank (MRR) reflects how early the correct answer appears in the ranked list by averaging the reciprocal of its first relevant position across all queries. Hit@k measures the proportion of queries for which at least one correct answer occurs within the top-$k$ retrieved results. Normalized Discounted Cumulative Gain (nDCG@k) assesses overall ranking quality by weighting relevant items according to their positions and normalizing by the ideal ranking, thereby capturing both accuracy and ordering of retrieved answers.
C.3 Implementation Details
We used EasyOCR [^2] https://github.com/JaidedAI/EasyOCR and Docling (Team, 2024) for detecting textual and visual elements, respectively. Whenever applicable, the answers are generated using a temperature of 0.0 to ensure deterministic results.
Family Model Parameters Knowledge Cutoff Release
GPT (OpenAI, 2025) gpt-4o \ Oct 2023 May 2024
Gemini (Anil et al., 2023) gemini-2.5-flash \ Jan 2025 June 2025
gemini-2.5-flash-lite \ Jan 2025 June 2025
gemini-2.0-flash \ June 2024 Jan 2025
Claude (Anthropic, 2025) claude-3-5-haiku-latest \ April 2024 Oct 2024
claude-opus-4-1-20250805 \ Jan 2025 Aug 2025
Llama (Grattafiori et al., 2024) Llama-3.2-11B-Vision-Instruct 11B \ Sep 2024
InternVL (Chen et al., 2024) InternVL3-8B 8B \ Apr 2025
Phi (Abdin et al., 2024) Phi-3-vision-128k-instruct 3.8B Oct 2023 July 2024
Qwen (Bai et al., 2025) Qwen2.5-VL-7B-Instruct 7B \ Feb. 2025
Qwen2.5-VL-32B-Instruct 32B \ Feb 2025
LLaVA (Liu et al., 2023) llava-1.5-7b-hf 7B Dec 2022 Apr. 2023
llava-1.5-13b-hf 13B Dec 2022 Apr. 2023
llava-v1.6-mistral-7b-hf 7B Dec 2023 Mar 2024
llava-v1.6-vicuna-13b-hf 13B Dec 2023 Mar 2024
Table 8: Overview of models used in the experiments, including their family, model name, parameter size, knowledge cutoff date, and release date.
Algorithm 1 Graph-based depth-first search for merging fragmented elements into coherent semantic units while preserving spatial layout.
Input: Set of OCR bounding boxes $B=\{b_{i}=(x_{1}^{i},y_{1}^{i},x_{2}^{i},y_{2}^{i},t_{i})\}_{i=1}^{|B|}$, distance threshold $\tau=15$
Output: Merged bounding boxes $\tilde{B}$
1:Initialize adjacency matrix $A\in\{0,1\}^{|B|\times|B|}$ with zeros
2:for $i=1$ to $|B|$ do
3:  for $j=i+1$ to $|B|$ do
4:   if $d_{\text{min}}(b_{i},b_{j})\leq\tau$ then
5:     $A[i,j]\leftarrow 1$, $A[j,i]\leftarrow 1$
6:   end if
7:  end for
8:end for
9:Initialize visited array $\text{visited}[1:|B|]\leftarrow\text{false}$
10:Initialize components list $\mathcal{C}\leftarrow[]$
11:for $i=1$ to $|B|$ do
12:  if $\text{visited}[i]=\text{false}$ then
13:   Initialize component $C\leftarrow[]$
14:   $\text{DFS}(i,A,\text{visited},C)$ $\triangleright$ Collect connected component
15:   Append $C$ to $\mathcal{C}$
16:  end if
17:end for
18:Initialize merged boxes $\tilde{B}\leftarrow$[]
19:for each component $C\in\mathcal{C}$ do
20:  if $|C|=1$ then
21:   Append $b_{C[0]}$ to $\tilde{B} \triangleright$ Single box, no merging
22:  else
23:   $x_{1}\leftarrow\min_{i\in C}x_{1}^{i}$, $y_{1}\leftarrow\min_{i\in C}y_{1}^{i}$
24:   $x_{2}\leftarrow\max_{i\in C}x_{2}^{i}$, $y_{2}\leftarrow\max_{i\in C}y_{2}^{i}$
25:   Sort $C$ by $(y_{1}^{i},x_{1}^{i})$ to get reading order
26:   $t_{\text{merged}}\leftarrow\text{" ".join}(\{t_{i}\}_{i\in\text{sorted}(C)})$
27:   Append $(x_{1},y_{1},x_{2},y_{2},t_{\text{merged}})$ to $\tilde{B}$
28:  end if
29:end for
30:return $\tilde{B}$
Appendix D Additional Experiments
D.1 Ablation Studies (Cont’d)
Element-level Reasoning Ensures Precision
Ablating the element agent causes a moderate yet consistent decline (–4.6 overall for GPT-4o; –6.3 for InternVL3-8B), especially numeric questions. Even with perfect retrieval, omitting fine-grained texts and layouts weakens factual grounding.
Impact of Global Thematic Guidance
Removing the global agent yields the smallest drop (–2.8 GPT-4o; –3.7 InternVL3-8B), as lower-level agents already embed partial global context via $\mathcal{K}_{p}$ and $\mathcal{K}_{e}$ in knowledge construction (Eq. 1/2). However, responses become less aware of overarching themes.
Subquery Generation Strengthens Retrieval.
Removing subqueries causes larger losses under retrieval (–5.0 GPT-4o; –11.3 InternVL3-8B) than with ground-truth pages (–2.9 to –4.9). Visual-aware subqueries markedly improve retriever accuracy by aligning textual intent with visual semantics–especially beneficial for weaker open-source encoders.
Appendix E Future Work
Element Parsing.
SlideAgent relies on advanced text and layout detection tools, which may struggle with visually complex or low-contrast designs (e.g., white charts on white backgrounds). Enhancing robustness in parsing could further improve overall reliability.
Modeling Element Relations.
SlideAgent currently treats elements independently. Extending it to explicitly model inter-element relations–for instance, via graph structures linking textual and visual components–offers a promising direction for capturing richer semantics, though at higher computational cost.
User-Centric Interaction.
While SlideAgent primarily enhances reasoning and retrieval for slide understanding, future work could make such agentic frameworks more user-centric through seamless integration into real-world user behaviors Wang et al. (2025b). Lightweight feedback mechanisms (e.g., contextual pop-ups or in-situ clarifications) could further improve interpretability and interaction without disrupting natural workflows or requiring excessive user input.
You are given a complete slide deck consisting of multiple slides. Your task is to synthesize a concise, high-level summary of the overall message, structure, and purpose of the deck.
### Cumulative Summary of the Preceding Slides {Cumulative Summary}
### Response format Please respond with a markdown string that follows the following format: Title <<Concise Explicit / Inferred Title of the Presentation>>
Objective <<What is the presentation trying to achieve? (e.g., inform, persuade, pitch, propose)>>
Structure Overview Slide 1: <<Brief description of the slide>> Slide 2: <<Brief description of the slide>> Slide N: <<Brief description of the slide>> Key Insights <<Major takeaway 1>> <<Major takeaway 2>> Audience <<Intended audience type (e.g., executives, investors, engineers)>>
Tone <<Overall tone: e.g., persuasive, analytical, optimistic, urgent>>
Table 9: Prompt template for the global agent to generate comprehensive slide deck summaries, including title, objective, structure, key insights, audience, and tone.
Element Type <<text | image | chart | table | icon | button | etc.>>
Position on Slide <<e.g., top-right, centered, below title>>
Verbatim Content <<if text, give the literal string; else describe the visual>>
Semantic Role <<What is the element trying to do or communicate?>>
Functional Purpose <<Its practical function within the slide (e.g., emphasize point, guide attention, show evidence, support action)>>
Relation to Slide <<How does it connect to or support the slide’s overall message?>>
Inferred Importance <<how central is this element to the slide? Answer with low, medium, or high>>
Table 10: Element-level slide description prompt format used by the element agent to generate detailed annotations for each visual component.
Model SlideVQA TechSlides FinSlides
Overall Num F1 Overall Num F1 Overall Num F1
Raw Models
Gemini 2.0 86.3 81.0 90.3 59.7 60.0 59.6 78.1 77.8 88.9
Gemini 2.5 89.0 85.7 93.2 61.5 65.0 59.5 76.6 76.4 83.3
Gemini 2.5-lite 81.8 75.5 90.1 56.3 55.0 57.0 78.1 78.4 66.7
Claude 4.1 85.7 82.4 89.7 58.0 77.5 48.3 52.6 52.0 60.8
Claude 3.5 58.2 64.0 50.9 55.8 83.7 42.5 47.8 48.0 40.3
GPT-4o 79.4 71.9 86.4 64.5 77.1 58.0 83.0 84.3 80.1
Multimodal RAG and Agentic Methods
ViDoRAG 81.8 73.8 87.0 65.8 78.0 58.7 84.2 84.7 81.5
SlideAgent 87.1 84.4 90.6 68.7 82.5 61.5 85.8 85.9 85.6
Impr. +7.7 +12.5 +4.2 +4.1 +5.4 +3.5 +2.8 +1.5 +5.5
Table 11: Performance comparison of baselines and proprietary models on SlideVQA, TechSlides, and FinSlides, assuming access to ground-truth pages containing the correct answer. All baseline methods use GPT-4o for question-answering.
Model SlideVQA TechSlides FinSlides
Overall Num F1 Overall Num F1 Overall Num F1
Raw Models
Llama 3.2 11B 44.6 52.1 34.6 47.1 62.8 39.3 39.1 39.2 33.5
Phi3 78.3 69.1 91.6 53.9 67.4 47.2 63.8 63.7 65.1
Qwen2.5 7B 85.1 77.7 95.3 57.7 69.8 51.8 52.7 52.0 77.8
Qwen2.5 32B 87.4 82.6 93.7 49.6 62.8 43.2 69.5 69.6 65.1
LLaVA 1.5 7B 42.9 27.9 79.9 24.6 15.0 39.4 14.2 14.4 20.9
LLaVA 1.5 13B 46.7 29.5 83.1 29.2 17.5 46.6 23.8 20.3 41.2
LLaVA 1.6 7B 59.0 45.8 84.2 36.2 40.0 34.0 16.0 15.2 21.3
LLaVA 1.6 13B 62.3 48.2 87.1 54.7 62.5 49.5 35.2 30.7 67.6
InternVL3 8B 73.3 65.4 85.7 58.4 72.1 51.5 56.3 55.9 65.6
Baseline methods based on InternVL3-8B
ViDoRAG 76.3 68.1 89.9 61.3 75.1 53.9 58.4 58.7 67.3
SlideAgent 82.8 75.3 93.3 64.6 79.5 58.0 62.8 62.6 68.1
Impr. +9.5 +9.8 +7.6 +6.2 +7.4 +6.4 +6.5 +6.7 +2.5
Table 12: Performance comparison of baselines and proprietary models on SlideVQA, TechSlides, and FinSlides, assuming access to ground-truth pages containing the correct answer. All baseline methods use InternVL3-8B for question-answering.

### v3

@@
 Multi-page visual documents such as manuals, brochures, presentations, and posters convey key information through layout, colors, icons, and cross-slide references.
-While large language models (LLMs) offer opportunities in document understanding, current systems struggle with complex, multi-page visual documents, particularly in fine-grained reasoning over elements and pages.
+While multimodal large language models (MLLMs) offer opportunities in document understanding, current systems struggle with complex, multi-page visual documents, particularly in fine-grained reasoning over elements and pages.
 We introduce SlideAgent, a versatile agentic framework for understanding multi-modal, multi-page, and multi-layout documents, especially slide decks.
@@
 During inference, SlideAgent selectively activates specialized agents for multi-level reasoning and integrates their outputs into coherent, context-aware answers.
-Extensive experiments show that SlideAgent achieves significant improvement over both proprietary (+7.9) and open-source models (+9.8).
+Extensive experiments show that SlideAgent significantly improves accuracy over both proprietary (+7.9%) and open-source models (+9.8%).
 SlideAgent: Hierarchical Agentic Framework for Multi-Page Visual Document Understanding
@@
 Visual documents–from earnings reports and academic lectures to business strategy presentations–are ubiquitous, conveying ideas not only from text, but also from the intricate interplay of layout, icons, visual hierarchy, and cross-page relationships.
-These documents are central to high-stakes domains such as finance, science, and technology.
+These documents are central to high-stakes domains such as finance, science, and technology, and small misunderstandings can cascade into costly decisions: analysts extract key metrics from earnings decks, engineers review tech specs to avoid implementation errors, and educators generate study materials whose inaccuracies can propagate at scale.
 Accurately interpreting them thus remains a pressing challenge.
-Figure 1: When given the full page, the LLM miscounts the number of product mix categories.
-After isolating the chart, it correctly identifies all eight categories, highlighting the importance of accurate element parsing.
+Figure 1: Accurate element parsing is important for fine-grained visual document understanding.
+The MLLM miscounts product-mix categories when reasoning over the entire slide, but succeeds after isolating the chart.
 Challenges.
@@
 State-of-the-art MLLMs process a limited number of images at once (Liu et al., 2023) and tend to treat each page holistically, missing fine-grained, element-level cues required for user queries (Faysse et al., 2024; Tanaka et al., 2025).
-In Figure 1, an LLM can miscount chart segments on a cluttered page of a slide deck, but correctly identifies all segments once the relevant chart is cropped–highlighting the importance of element-level parsing for latent reasoning abilities.
+In Figure 1, an MLLM can miscount chart segments on a cluttered page of a slide deck, but correctly identifies all segments once the relevant chart is cropped–highlighting the importance of element-level parsing for latent reasoning abilities.
 2) Domain-Specific Visual Semantics.
@@
 Icons convey abstract concepts (lightbulbs for innovation, arrows for causal relationships); and spatial positioning signals importance (centered elements typically matter more than corner annotations).
-falter in spatial reasoning Wang et al.
+However, current MLLMs falter in spatial reasoning Wang et al.
 (2024b, a), failing to locate visual elements Sharma and Vats (2025); Bhattacharyya et al.
@@
 This Work.
-We present SlideAgent, an LLM-based agentic framework for fine-grained understanding of multi-page, varying-size visual documents.
+We present SlideAgent, an MLLM-based agentic framework for fine-grained understanding of multi-page, varying-size visual documents.
 Inspired by the human information processing model (Lang, 2000; Naysmith et al., 2021), SlideAgent employs a hierarchical architecture with specialized agents at three levels: global (document-wide topics), page (page-specific features and cross-page relations), and element (fine-grained components such as charts, figures, and text blocks).
@@
 We benchmark SlideAgent and baseline models, demonstrating significant performance on both open-source and proprietary models.
-SlideAgent surpasses its LLM counterpart by 7.9 (proprietary) and 9.8 (open-source), respectively.
+SlideAgent surpasses its MLLM counterpart by 7.9% (proprietary) and 9.8% (open-source) in accuracy, respectively.
 We also show that SlideAgent enhances spatial reasoning, visual counting, and cross-element understanding, with results that are highly interpretable.
@@
 At each level, specialized agents generate query-agnostic knowledge during knowledge construction, then retrieve and reason over query-specific textual & visual knowledge during inference stage.
-Sample knowledge $\mathcal{K}$ generated by SlideAgent is in Appendix Figure 7,8,9 and answers generated by the agents are in Figure 5.
+Sample knowledge $\mathcal{K}$ generated by SlideAgent is in Appendix Figure 10,11,12 and answers generated by the agents are in Figure 8.
 2.1 Knowledge Construction Stage
@@
 Global Agent
-The global agent $\mathcal{M}_{g}$ generates document-level knowledge $\mathcal{K}_{g}=\mathcal{M}_{g}(\mathcal{P})$, capturing the overall summary, objectives, and narrative flow of the document.
+The global agent $\mathcal{M}_{g}$ generates initial document-level knowledge $\mathcal{K}_{g}^{(0)}$, capturing the overall summary, objectives, and narrative flow of the document.
 This layer establishes overarching themes to support high-level reasoning about the document’s purpose.
-Since visual documents are often large and LLMs have a limited capacity for processing visuals, SlideAgent samples the first three pages to generate $\mathcal{K}_{g}$ (example in Appendix Figure 7).
+Since visual documents are often large and MLLMs have a limited capacity for processing visuals, SlideAgent samples the first three pages to generate $\mathcal{K}_{g}^{(0)}$ (Appendix Figure 10).
 Page Agent
-For each page $p_{i}\in\mathcal{P}$, the page agent generates page-level knowledge $\mathcal{K}_{p}$ in a sequential manner, conditioned on the page’s visual content $v_{i}$, the global knowledge $\mathcal{K}_{g}$, and the knowledge from the preceding page $\mathcal{K}_{p}^{i-1}$:
-$\mathcal{K}_{p}^{i}=\mathcal{M}_{p}(v_{i},\mathcal{K}_{g},\mathcal{K}_{p}^{i-1}),i\in[1,|\mathcal{P}|].\quad\mathcal{K}_{p}^{0}=\emptyset$ (1)
+For each page $p_{i}\in\mathcal{P}$, the page agent generates page-level knowledge $\mathcal{K}_{p}$ in a sequential manner, conditioned on the page’s visual content $v_{i}$, the initial global knowledge $\mathcal{K}_{g}^{(0)}$, and the knowledge from the preceding page $\mathcal{K}_{p}^{i-1}$:
+$\mathcal{K}_{p}^{i}=\mathcal{M}_{p}(v_{i},\mathcal{K}_{g}^{(0)},\mathcal{K}_{p}^{i-1}),i\in[1,|\mathcal{P}|].\mathcal{K}_{p}^{0}=\emptyset$ (1)
 The complete page-level knowledge $\mathcal{K}_{p}=\bigcup_{i=1}^{|\mathcal{P}|}\mathcal{K}_{p}^{i}$ provides an intermediate representation that captures page-specific content while linking them to the global context (sample in Figure 11).
-To ensure comprehensive understanding of all pages, $\mathcal{K}_{p}$ is subsequently used to refine $\mathcal{K}_{g}$.
+To ensure comprehensive understanding of all pages, $\mathcal{K}_{p}$ is subsequently used to refine $\mathcal{K}_{g}^{(0)}$ through a single-pass fieldwise rewrite:
+$\mathcal{K}_{g}=\mathcal{M}_{g}^{\mathrm{refine}}(\mathcal{K}_{g}^{(0)},\mathrm{concat}(\mathcal{K}_{p}^{1},\ldots,\mathcal{K}_{p}^{|\mathcal{P}|})).$ (2)
+The global agent regenerates each global field from the complete page-level evidence, using $\mathcal{K}_{g}^{(0)}$ only as a weak prior for terminology and early context.
+This rewrite can overwrite earlier hypotheses when later pages provide stronger evidence, reducing bias toward the first pages.
+The refinement prompt is shown in Appendix Table 13.
 Element Agent
-LLMs often struggle with spatial reasoning over visual documents Wang et al.
+MLLMs often struggle with spatial reasoning over visual documents Wang et al.
 (2024b, a), failing to locate visual elements Sharma and Vats (2025); Bhattacharyya et al.
@@
 To address this, our element agent integrates external tools to explicitly capture the spatial and structural information of each page.
+Case Example Agents
+Global Understanding Asks about the overall theme, purpose, or general summary of the entire presentation. "What is the main topic of the presentation?", "What is this deck about?" Global
+Fact-based Direct Query. Asks for specific facts, data, or information from particular slides. "What is the revenue reported on slide 7?", "Which slide shows the product roadmap?" Page, Element
+Multi-hop Reasoning Requires comparing information across multiple slides or elements. "Compare revenues in slide 5 and slide 10", "How do Sweden and Denmark compare?" Global, Page, Element
+Layout / Visual Relationship Asks about visual relationships, positioning, or layout elements.
+"What does the diagram below the table on slide 12 illustrate?", "Is the color red used to denote negative performance?" Element
+Uncertain If the query is unclear, use all agents to answer the query. \ Global, Page, Element
+Table 1: SlideAgent’s query classification that determines which hierarchical agents to activate.
 At the finest granularity, the element agent decomposes each page $p_{i}$ into a set of elements using a layout parsing pipeline $f:v_{i}\rightarrow\{(i,e_{j},b_{j},t_{j})\}_{j=1}^{M_{i}}$, where each element is represented by its page index $i$, verbatim text $e_{j}$, bounding box coordinates $b_{j}$, and element type $t_{j}$.
-$f$ integrates text detection, layout detection, and element classification, followed by post-processing to merge fragmented elements (Appendix A.1).
+$f$ integrates text detection, layout detection, and element classification, followed by post-processing to merge fragmented elements (Appendix B.1).
 For each detected element $e_{j}$, the agent consumes the annotated visuals and metadata to generate element-level knowledge:
@@
 Thus, the agent orchestrator first attempts to classify each query $q$ into one of 4 predefined categories, such as global understanding, fact-based direct queries, multi-hop reasoning, and layout & visual relationships.
-Each type corresponds to a question-answering strategy and a targeted set of agents (Appendix Table 7).
+Each type corresponds to a question-answering strategy and a targeted set of agents (Table 1).
 For instance, global understanding queries activate only the global agent, as they require a broad overview, while fact-based queries trigger both the page and element agents to retrieve detailed, slide-specific information.
@@
 Multimodal LLMs (Type 1)
-Gemini 2.0 75.0 71.3 79.8 50.4 67.6 41.6 70.8 70.6 77.8
-Gemini 2.5 83.8 78.3 91.8 51.1 71.4 41.2 76.2 75.8 100.0
-Gemini 2.5-lite 71.2 60.8 87.0 47.3 58.1 41.9 57.0 56.6 68.3
-Claude 4.1 78.4 74.3 82.3 61.0 81.4 52.3 56.5 54.8 73.3
-Claude 3.5 62.5 68.3 54.6 52.5 80.2 39.5 48.5 49.5 29.6
+Gemini-2.0-Flash 75.0 71.3 79.8 50.4 67.6 41.6 70.8 70.6 77.8
+Gemini-2.5-Flash 83.8 78.3 91.8 51.1 71.4 41.2 76.2 75.8 100.0
+Gemini-2.5-Flash-Lite 71.2 60.8 87.0 47.3 58.1 41.9 57.0 56.6 68.3
+Claude-4.1-Opus 78.4 74.3 82.3 61.0 81.4 52.3 56.5 54.8 73.3
+Claude-3.5-Haiku 62.5 68.3 54.6 52.5 80.2 39.5 48.5 49.5 29.6
 GPT-4o 77.0 72.1 84.0 63.4 78.3 53.9 80.0 80.8 62.1
-Multimodal RAG (Type 2) and Agentic Methods (Type 3)
+Multimodal RAG (Type 2) and Agentic Methods (Type 3) with GPT-4o as base models.
 COLPALI 78.8 73.7 83.4 64.1 73.2 54.5 80.9 81.5 62.7
@@
 The best and second best performance are highlighted in bold and underlined.
-SlideAgent outperforms all baseline methods (Type 2/3) sharing the same base model (GPT-4o), and even outperforms stronger raw LLMs (e.g.
+SlideAgent outperforms all baseline methods (Type 2/3) sharing the same base model, and even outperforms stronger raw LLMs (e.g.
 Gemini-2.5).
+late
 Datasets
@@
 Models.
-We benchmark SlideAgent against 3 types of baselines: 1) Multimodal LLMs: 15 LLMs from 8 model families, including proprietary models (GPT-4o (OpenAI, 2025), Gemini (Anil et al., 2023), Claude (Anthropic, 2025)) and open-source models (Llama-3.2 (Grattafiori et al., 2024), InternVL3 (Chen et al., 2024), Phi-3 (Abdin et al., 2024), Qwen2.5-VL (Bai et al., 2025)); 2) Multimodal RAG Methods: VisRAG Yu et al.
+We benchmark SlideAgent against 3 types of baselines: 1) Multimodal LLMs: 15 MLLMs from 8 model families, including proprietary models (GPT-4o (OpenAI, 2025), Gemini (Anil et al., 2023), Claude (Anthropic, 2025)) and open-source models (Llama-3.2 (Grattafiori et al., 2024), InternVL3 (Chen et al., 2024), Phi-3 (Abdin et al., 2024), Qwen2.5-VL (Bai et al., 2025)); 2) Multimodal RAG Methods: VisRAG Yu et al.
 (2025), VDocRAG Tanaka et al.
@@
 (2024); 3) Multi-agent Systems: ViDoRAG Wang et al.
-(2025a).
-We evaluate SlideAgent using two backbone LLMs, including both proprietary models (GPT-4o (OpenAI, 2025)) and open-source models (InternVL3-8B (Chen et al., 2024)).
+(2025a) and MDocAgent (Han et al., 2025).
+We evaluate SlideAgent using two backbone MLLMs, including both proprietary models (GPT-4o (OpenAI, 2025)) and open-source models (InternVL3-8B (Chen et al., 2024)).
 These models are chosen for their widespread use in state-of-the-art QA systems Yu et al.
@@
 Otherwise, each model receives as many retrieved pages up to its input capacity.
-Figure 3: Performance comparison among variants of SlideAgent with base model GPT-4o.
 Metrics.
@@
 Model Overall Num F1
-Raw LLMs
-Gemini 2.0 72.3 66.1 81.3
-Gemini 2.5 86.9 81.1 95.4
-Gemini 2.5 lite 71.7 62.6 87.0
-Claude 4.1 36.3 35.4 38.9
-Claude 3.5 31.5 30.8 37.0
+Multimodal LLMs
+Gemini-2.0-Flash 72.3 66.1 81.3
+Gemini-2.5-Flash 86.9 81.1 95.4
+Gemini-2.5-Flash Lite 71.7 62.6 87.0
+Claude-4.1-Opus 36.3 35.4 38.9
+Claude-3.5-Haiku 31.5 30.8 37.0
 GPT-4o 69.0 59.3 90.5
@@
 Impr. +10.6 +10.5 +3.6
-Table 2: Performance comparison among Gemini, Claude, GPT-4o, ViDoRAG, and SlideAgent models.
+Table 3: Performance comparison among proprietary models, such as Gemini, Claude, GPT-4o, ViDoRAG, and SlideAgent on InfoVQA Mathew et al. (2022).
 3.1 End-to-End Performance
-Tables 1/2/3 and Appendix Table 5 demonstrate the end-to-end performance of SlideAgent across proprietary and open-source models.
+Tables 2/3/4 and Appendix Table 5 demonstrate the end-to-end performance of SlideAgent across proprietary and open-source models.
 Model SlideVQA TechSlides FinSlides
@@
 Multimodal LLMs (Type 1)
-Llama 3.2 11B 42.9 43.3 42.3 41.4 52.5 36.2 23.3 23.2 26.2
-Phi3 72.3 61.8 90.6 59.4 60.0 59.1 48.8 48.5 64.3
-Qwen2.5 7B 79.5 70.5 94.3 59.3 52.5 65.9 53.6 52.5 85.7
-Qwen2.5 32B 79.2 71.1 92.2 67.5 87.5 60.6 57.4 56.6 87.5
-LLaVA 1.5 7B 36.8 22.4 79.0 23.3 12.5 38.7 10.7 10.1 16.6
-LLaVA 1.5 13B 44.9 25.1 81.8 28.1 17.5 45.0 20.6 16.5 36.7
-LLaVA 1.6 7B 50.9 37.3 82.6 34.4 37.5 32.2 12.2 12.1 17.8
-LLaVA 1.6 13B 16.7 10.2 81.5 45.2 40.0 49.1 32.0 31.3 64.3
-InternVL3 8B 63.0 56.5 74.1 55.4 57.5 54.4 49.8 49.5 64.3
-Multimodal RAG (Type 2) and Agentic Method (Type 3)
+Llama-3.2-11B-Vision-Instruct 42.9 43.3 42.3 41.4 52.5 36.2 23.3 23.2 26.2
+Phi-3-vision-128k-instruct 72.3 61.8 90.6 59.4 60.0 59.1 48.8 48.5 64.3
+Qwen2.5-VL-7B-Instruct 79.5 70.5 94.3 59.3 52.5 65.9 53.6 52.5 85.7
+Qwen2.5-VL-32B-Instruct 79.2 71.1 92.2 67.5 87.5 60.6 57.4 56.6 87.5
+llava-1.5-7b-hf 36.8 22.4 79.0 23.3 12.5 38.7 10.7 10.1 16.6
+llava-1.5-13b-hf 44.9 25.1 81.8 28.1 17.5 45.0 20.6 16.5 36.7
+llava-v1.6-mistral-7b-hf 50.9 37.3 82.6 34.4 37.5 32.2 12.2 12.1 17.8
+llava-v1.6-vicuna-13b-hf 16.7 10.2 81.5 45.2 40.0 49.1 32.0 31.3 64.3
+InternVL3-8B 63.0 56.5 74.1 55.4 57.5 54.4 49.8 49.5 64.3
+Multimodal RAG (Type 2) and Agentic Method (Type 3) with InternVL3-8B as base models.
 COLPALI 63.4 56.7 73.8 57.1 60.9 55.2 50.4 49.3 65.7
@@
 Impr. +9.8 +11.7 +5.4 +7.7 +20.5 +2.3 +13.5 +13.3 +4.0
-Table 3: Performance comparison of various models on SlideVQA, TechSlides, and FinSlides datasets. SlideAgent outperforms all baseline methods (Type 2/3) sharing the same base model (GPT-4o).
+Table 4: Performance (correctness %) of various models on SlideVQA, TechSlides, and FinSlides dataset. SlideAgent outperforms all baseline methods (Type 2/3) sharing the same LLM backbone.
 Consistent Improvements across Architectures
 For both proprietary and open-source models, SlideAgent consistently outperforms all baseline methods (Type 2/3) and the base models across all metrics.
-For proprietary models (Table 1), SlideAgent improves overall accuracy by +7.9, numeric reasoning by +8.3, and lexical overlap by +6.5 on SlideVQA.
-Similar trends are observed for TechSlides (+7.5 overall; +12.3 F1) and FinSlides (+5.5 overall; +17.5 F1), indicating robust performance across diverse domains.
-For open-source models, the advantage remains pronounced: SlideAgent outperforms InternVL3-8B by +9.8 overall (+11.7 numeric), showing that our hierarchical, multi-agent design generalizes well across LLMs.
+For proprietary models (Table 2), SlideAgent improves overall accuracy by +7.9+7.9%, numeric reasoning by +8.3+8.3%, and lexical overlap by +6.5+6.5% on SlideVQA.
+Similar trends are observed for TechSlides (+7.5+7.5% overall; +12.3+12.3% F1) and FinSlides (+5.5+5.5% overall; +17.5+17.5% F1), indicating robust performance across diverse domains.
+For open-source models, the advantage remains pronounced: SlideAgent outperforms the original MLLM by +9.8+9.8% overall (+11.7+11.7% numeric), showing that our hierarchical, multi-agent design generalizes well across MLLMs.
+Figure 3: Absolute performance gains over MDocAgent Han et al.
+(2025) when using GPT-4o and InternVL3-8B as the backbone models on SlideVQA Tanaka et al.
+(2023).
+SlideAgent yields significant overall improvements across both backbones.
 Comparison with Multimodal LLMs
-SlideAgent consistently achieves the best or second-best performance among proprietary models (Tables 1 and  11).
-Notably, although the base model slightly lags behind stronger LLMs such as Gemini 2.5 in raw capability (e.g.
-83.8 for Gemini 2.5 vs.
-77.0 for GPT-4o in Table 1), SlideAgent’s structured reasoning pipeline fills this gap (84.9 overall for SlideAgent).
-For open-source models, SlideAgent achieves better performance across all models except for the Qwen2.5 family.
-While strong base models such as Gemini-2.5 and Qwen2.5 already exhibit advanced multimodal comprehension, SlideAgent is model-agnostic in nature and can be directly applied to these models to further enhance performance.
+SlideAgent consistently achieves the best or second-best performance among proprietary models (Table 2 and Appendix Table 11).
+Notably, although the base model slightly lags behind stronger MLLMs such as Gemini-2.5-Flash in raw capability (e.g.
+83.8 for Gemini-2.5-Flash vs.
+77.077.0% for GPT-4o in Table 2), SlideAgent’s structured reasoning pipeline fills this gap (84.984.9% overall for SlideAgent).
+For open-source models, SlideAgent achieves better performance across all models except for the Qwen2.5-VL family.
+While strong base models such as Gemini-2.5 and Qwen2.5-VL already exhibit advanced multimodal comprehension, SlideAgent is model-agnostic in nature and can be directly applied to these models to further enhance performance.
 Figure 4: Accuracy of SlideAgent and base models (GPT-4o / InternVL3-8B) on different query types.
@@
 SlideAgent improves question-answering across diverse query types, especially in multi-hop reasoning and visual/layout questions.
-The greatest improvement occurs on case 3 (multi-hop reasoning), with a 9.8 point improvement (67.4 to 77.2).
+The greatest improvement occurs on case 3 (multi-hop reasoning), with a 9.89.8% improvement (67.467.4% to 77.277.2%).
 This means explicitly guiding the model’s reasoning using generated knowledge $(\mathcal{K}_{g},\mathcal{K}_{p},\mathcal{K}_{e})$ significantly improves reasoning capabilities.
-A notable 7.7-point improvement (66.7 $\rightarrow$ 74.4) is also observed in visual/layout reasoning, demonstrating the benefit of fine-grained element-level reasoning and retrieval that multimodal LLMs hardly achieve.
-Both SlideAgent and its counterpart perform well on Case 2 (Fact-based Direct Queries), with only a modest 2.1-point improvement, reflecting the model’s ability to handle page-level reasoning with little space for further enhancement.
+A notable 7.77.7% improvement (66.766.7% $\rightarrow 74.4$%) is also observed in visual/layout reasoning, demonstrating the benefit of fine-grained element-level reasoning and retrieval that multimodal LLMs hardly achieve.
+Both SlideAgent and its counterpart perform well on Case 2 (Fact-based Direct Queries), with only a modest 2.12.1% improvement, reflecting the model’s ability to handle page-level reasoning with little space for further enhancement.
 3.2 Performance with Ground-truth Pages
@@
 As all models receive the exact pages with the answers, noise introduced in retrieval is eliminated.
-Under this oracle setting, SlideAgent still improves over the base model (+7.7 overall and +12.5 numeric on SlideVQA), demonstrating the effectiveness of element-level retrieval.
-Text-based Retrievers MRR Recall@1 nDCG@1 Recall@3 Hit@3 nDCG@3
-BM25 Robertson et al. (2004) 59.0 51.5 52.0 63.0 65.3 57.7
-w/ SA 63.9 +4.9 54.9 +3.4 56.6 +4.6 67.1 +4.1 68.8 +3.5 62.6 +4.9
-BGE Xiao et al. (2023) 70.1 56.1 60.9 75.8 78.1 69.2
-w/ SA 72.3 +2.2 58.4 +2.3 63.2 +2.3 77.4 +1.6 80.3 +2.2 71.3 +2.1
-SFR Meng et al. (2024) 70.1 56.1 60.9 75.8 78.1 69.2
-w/ SA 76.5 +6.4 59.9 +3.8 69.4 +8.5 77.3 +1.5 81.8 +3.7 73.2 +4.0
-Multimodal Retrievers MRR Recall@1 NDCG@1 Recall@3 Hit@3 NDCG@3
-SigLIP2 Tschannen et al. (2025) 26.7 15.9 18.0 31.0 34.0 25.0
-w/ SA 28.0 +1.3 16.3 +0.4 18.0 +0.0 32.3 +1.3 35.5 +1.5 26.1 +1.1
-COLPALI Faysse et al. (2024) 82.1 68.2 75.5 78.9 84.0 77.4
-w/ SA 82.9 +0.8 70.4 +2.2 76.2 +0.7 88.6 +9.7 90.1 +6.1 83.1 +5.7
-VisRAG Yu et al. (2025) 76.0 63.3 68.6 82.3 84.1 76.0
-w/ SA 79.7 +3.7 66.3 +3.0 71.6 +3.0 85.5 +3.2 87.7 +3.6 79.4 +3.4
-Table 4: Artifacts generated by SlideAgent, particularly $\hat{Q}$ and $\mathcal{M}_{p}$, improves page-level retrieval performance, especially for text-based retrievers.
+Under this oracle setting, SlideAgent still improves over the base model (+7.7+7.7% overall and +12.5+12.5% numeric on SlideVQA), demonstrating the effectiveness of element-level retrieval.
+3.3 Error Analysis
+We annotate 40 SlideVQA failure cases to quantify parsing-related error propagation.
+As shown in Table 5, only OCR Error and Tiny Visuals are directly attributable to parsing failures, accounting for 5/40 cases (12.5%).
+The dominant categories are Ambiguous Question (22.5%), Answer Mislocation (17.5%), and Valid Alternative Answer (15.0%), indicating that ambiguity, retrieval drift, and annotation subjectivity account for a larger share of failures than OCR/layout parsing.
+Figure 5: Error analysis of 40 failure cases on SlideVQA. Only OCR Error and Tiny Visuals (12.5% combined) are directly attributable to parsing failures.
 3.4 Effectiveness of Knowledge Construction
@@
 Specifically, we test page-level retrieval using generated subqueries $\hat{Q}$ and page knowledge $\mathcal{K}_{p}$ from SlideAgent.
-Table 4 shows consistent gains across both text-based retrievers (BM25 Robertson et al.
+Figure 6 shows significant gains across both text-based retrievers (BM25 Robertson et al.
 (2004), BGE Xiao et al.
@@
 (2025)).
+Figure 6: Artifacts generated by SlideAgent, particularly $\hat{Q}$ and $\mathcal{M}_{p}$, improves page-level retrieval performance, especially for text-based retrievers.
 Text-based Retrievers Show Largest Gains.
@@
 VisRAG Yu et al.
-(2025) achieves a larger ranking boost (+3.7 MRR), suggesting that multimodal LLMs still benefit from textual guidance in aligning queries to page content.
-SigLIP2 shows minimal gains (+1.3 MRR), likely because it is optimized for natural image domains and transfers less effectively to document-style inputs.
-Figure 5: Example answers generated by SlideAgent. Agents at different levels work together to provide comprehensive responses.
+(2025) achieves a larger ranking boost (+3.7+3.7% MRR), suggesting that multimodal LLMs still benefit from textual guidance in aligning queries to page content.
+SigLIP2 shows minimal gains (+1.3+1.3% MRR), likely because it is optimized for natural image domains and transfers less effectively to document-style inputs.
+Figure 7: Performance comparison among variants of SlideAgent with base model GPT-4o.
 3.5 Ablation Studies
 We analyze each design choice by removing global agent (w/o G), page agent (w/o P), element agent (w/o E), and subquery generation (w/o S).
-Figures 3/6 summarize results across proprietary (GPT-4o) and open-source (InternVL3-8B) backbones.
+Figures 7 summarize results across proprietary (GPT-4o) and open-source (InternVL3-8B) backbones.
 Page-level Reasoning is Critical
@@
 Its absence breaks this structural bridge between global context and page details.
-Results for SlideAgent w/o E, w/o G, and w/o S are in Appendix D.1.
+Element-level Reasoning Ensures Precision
+Ablating the element agent causes a moderate yet consistent decline (–4.6 overall for GPT-4o; –6.3 for InternVL3-8B), especially numeric questions.
+Even with perfect retrieval, omitting fine-grained texts and layouts weakens factual grounding.
+Impact of Global Thematic Guidance
+Removing the global agent yields the smallest drop (-2.82.8% GPT-4o; –$3.7$ InternVL3-8B), as lower-level agents already embed partial global context via $\mathcal{K}_{p}$ and $\mathcal{K}_{e}$ in knowledge construction (Eq.
+1/3).
+However, responses become less aware of overarching themes.
+Subquery Generation Strengthens Retrieval.
+Removing subqueries causes larger losses under retrieval (–5.0 GPT-4o; –11.3 InternVL3-8B) than with ground-truth pages (–2.9 to –4.9).
+Visual-aware subqueries markedly improve retriever accuracy by aligning textual intent with visual semantics–especially beneficial for weaker open-source encoders.
 3.6 Qualitative Answer Analysis
 Figure 8 shows a qualitative example of answer generated by SlideAgent to illustrate how SlideAgent leverages hierarchical knowledge for reasoning.
+Figure 8: Example answers by SlideAgent. Agents at different levels collaborate for comprehensive responses.
 Given the slide deck, SlideAgent builds hierarchical knowledge $\mathcal{K}=\{\mathcal{K}_{g},\mathcal{K}_{p},\mathcal{K}_{e}\}$ (Figures 1012).
@@
 4 Related Work
-In this section, we discuss relevant works in visual document understanding and Multimodal LLMs. A more detailed description is in Appendix B.
+In this section, we discuss relevant works in visual document understanding and Multimodal LLMs. A more detailed description is in Appendix A.
 Visual Document Understanding relied on computer vision pipelines combining OCR, layout parsing, and heuristics to extract semantics Shilman et al.
@@
 (2024), these models struggle with visual documents such as slides, which require precise grounding of heterogeneous elements (charts, tables, icons).
-5 Conclusion
-We present SlideAgent, a hierarchical agentic framework that leverages specialized agents at multiple levels for fine-grained infographics understanding, particularly for slide decks.
-SlideAgent demonstrates significant performance gains and strong generalizability across both proprietary and open-source models.
+5 Conclusion & Future Work
+We present SlideAgent, a hierarchical agentic framework that leverages specialized agents at multiple granularity levels for multi-page visual document understanding, particularly for slide decks.
+Future work can improve parsing robustness on low-contrast layouts, model inter-element relations via graphs, replace the first-$N$-pages heuristic with content-aware page selection for $\mathcal{K}_{g}$, and extend SlideAgent to multi-turn scenarios.
 6 Limitations
-We did not evaluate element-level retrieval due to the lack of corresponding annotations in existing datasets and the difficulty of defining precise boundaries among elements.
-Benchmarks in visual document understanding could include such annotations to enable finer-grained analysis.
-Retrieval Method.
-Our framework supports both textual and multimodal retrievers.
-We primarily adopt text-based retrieval for efficiency, though future work could explore multimodal retrievers or domain-specific strategies tailored to slide decks.
-7 Ethical Considerations
-Data Privacy, Consent, and Intellectual Property.
-Visual documents such as those processed by SlideAgent may contain sensitive business or personal data.
-Ensuring compliance with privacy regulations (e.g., GDPR, CCPA) and obtaining appropriate consent are essential.
-Organizations should also respect intellectual property rights and establish policies that balance knowledge sharing with fair use.
-Content Reliability and User Responsibility.
-Like other LLM-based systems, SlideAgent may inherit biases or produce inaccurate content.
-While it enhances question answering, outputs should be verified by human judgment, particularly in sensitive or high-stakes scenarios.
-Clear user guidance and validation practices can help ensure responsible use.
+1) Element-level Retrieval.
+We assess element-level grounding on a manually annotated subset of SlideVQA questions, providing direct evidence of fine-grained retrieval.
+However, element boundaries may vary across OCR and layout-parsing tools.
+Broader benchmarks with element-level annotations would enable more comprehensive evaluation.
+2) Retrieval Method.
+Our framework supports both textual and multimodal retrieval.
+For efficiency, we primarily adopt text-based retrieval.
+Future work could explore multimodal retrievers or domain-specific retrieval strategies.
 Disclaimer
@@
 This document is not intended as investment research or investment advice, or a recommendation, offer or solicitation for the purchase or sale of any security, financial instrument, financial product or service, or to be used in any way for evaluating the merits of participating in any transaction, and shall not constitute a solicitation under any jurisdiction or to any person, if such solicitation under such jurisdiction or to such person would be unlawful.
-Appendix A Details Method
-A.1 Element Merging
-Raw output, especially text spans from the element processing pipeline (Section 2.1) can fragment coherent texts into multiple parts, creating challenges for downstream understanding.
-We thus adopt a graph-based merging algorithm to reconstruct semantically coherent text blocks while preserving spatial layout.
-Distance-Based Adjacency
-Our merging criterion is based on minimum distance between bounding boxes.
-For two boxes $b_{i}=(x_{1}^{i},y_{1}^{i},x_{2}^{i},y_{2}^{i})$ and $b_{j}=(x_{1}^{j},y_{1}^{j},x_{2}^{j},y_{2}^{j})$, the minimum distance is computed from the horizontal and vertical distances $d_{h}$ and $d_{v}$:
-$d_{\text{min}}(b_{i},b_{j})=\sqrt{d_{h}^{2}+d_{v}^{2}}$
-where $d_{h}=\min(|x_{2}^{i}-x_{1}^{j}|,|x_{2}^{j}-x_{1}^{i}|)$ if boxes don’t overlap horizontally, and $d_{h}=0$ otherwise.
-Similarly, $d_{v}=\min(|y_{2}^{i}-y_{1}^{j}|,|y_{2}^{j}-y_{1}^{i}|)$ if boxes don’t overlap vertically, and $d_{v}=0$ otherwise.
-Two boxes are considered adjacent if $d_{\text{min}}(b_{i},b_{j})\leq\tau$ where $\tau$ is a threshold we choose to be 15 pixels.
-The results are fed into the following graph-based component detection.
-Notation Description
-$p_{i},v_{i}$ A slide page and its visual content
-$\mathcal{P}=\{p_{1},p_{2},\dots\}$ Set of pages in the slide deck
-$q$ User query
-$\hat{Q}$ Set of subqueries generated from the original query $q$
-$a$ The final answer generated by SlideAgent
-$e_{j}$ Individual element in a slide, consisting of visual content, bounding box, and type
-$b_{j}$ Bounding box coordinates of an element within a slide
-$t_{j}$ Type of the element (e.g., text, chart, image, table)
-$\mathcal{K}_{g},\mathcal{K}_{p},\mathcal{K}_{e}$ Global, page, and element knowledge
-$\mathcal{M}_{g},\mathcal{M}_{p},\mathcal{M}_{e}$ Global, page, and element agent
-$\mathcal{R}$ Retrieval function that fetches relevant pages or elements based on subqueries
-$\phi(\cdot)$ Answer synthesizer that combines reasoning from all agents to generate the final answer $a$
-[…]
-$h_{g},h_{p},h_{e}$ Answers and reasoning from global, page, and element agents
-Table 6: Mathematical notation used throughout the paper.
-Case Characteristics Example Agents
-1-Global Understanding Asks about the overall theme, purpose, or general summary of the entire presentation. "What is the main topic of the presentation?", "What is this deck about?" Global
-2-Fact-based Direct Query Asks for specific facts, data, or information from particular slides. "What is the revenue reported on slide 7?", "Which slide shows the product roadmap?" Page, Element
-3-Multi-hop Reasoning Requires comparing information across multiple slides or elements. "Compare revenues in slide 5 and slide 10", "How do Sweden and Denmark compare?" Global, Page, Element
-4-Layout / Visual Relationship Asks about visual relationships, positioning, or layout elements.
-"What does the diagram below the table on slide 12 illustrate?", "Is the color red used to denote negative performance?" Element
-5-Uncertain If the query is unclear, use all agents to answer the query. \ Global, Page, Element
-Table 7: SlideAgent’s query classification that determines which hierarchical agents to activate.
-Appendix B Extended Related Work
-B.1 Document and Infographics Understanding
+Acknowledgements
+The authors would like to thank Nishan Srishankar, Kelly Patel, and Manuela Veloso at J.P. Morgan AI Research for their valuable discussions and feedback on this work.
+Appendix A Extended Related Work
+A.1 Document and Infographics Understanding
 Early research in document and infographic understanding predates multimodal LLMs and was dominated by computer vision pipelines.
@@
 These advances have achieved impressive results on single-image document VQA tasks by unifying textual, layout, and visual signals, laying the foundation for more recent LLM-based systems.
-B.1.1 Slide Deck Understanding
+A.1.1 Slide Deck Understanding
 Building on document and infographic understanding, research has also explored presentation slides, which pose unique challenges due to their combination of dense text, figures, and multi-page structure.
@@
 (2024) have emphasized long-context processing over lengthy multi-page PDFs, highlighting the scalability issues that arise when moving from single documents to multi-slide collections.
-B.1.2 Slide Visual Question Answering
+A.1.2 Slide Visual Question Answering
 Slide visual question answering (Slide VQA) emerges as a prominent direction within this broader area, aiming to automatically interpret presentation slides to answer natural-language queries.
@@
 Collectively, these datasets and methods underscore both the promise and complexity of slide understanding, while leaving open the question of whether element-level reasoning—beyond page-level retrieval—can substantially improve performance.
-B.1.3 General Purpose LLMs
+A.1.3 General Purpose LLMs
 General-purpose large language models (MLLMs) such as GPT-4 Achiam et al.
@@
 Visual prompting methods like Set-of-Marks Yang et al.
-(2023) further enhance reasoning by augmenting inputs with structured visual annotations.
+(2023) enhance reasoning by augmenting inputs with structured visual annotations.
+Multi-agent systems Wang et al.
+(2025b, 2026) have also enhanced the multi-perspective reasoning capabilities of LLM-powered systems.
 While these open-source and closed-source MLLMs form the foundation for many comparisons, they often struggle when applied to slides, where precise grounding and fine-grained interpretation are required to parse heterogeneous elements such as charts, tables, and icons.
@@
 This design enables robust comprehension across diverse slide domains and opens the door to systematic evaluation of whether element-level reasoning provides measurable improvements over page-level retrieval alone.
+Appendix B Additional Method Details
+B.1 Element Merging
+Raw output, especially text spans from the element processing pipeline (Section 2.1) can fragment coherent texts into multiple parts, creating challenges for downstream understanding.
+We thus adopt a graph-based merging algorithm to reconstruct semantically coherent text blocks while preserving spatial layout.
+Distance-Based Adjacency
+Our merging criterion is based on minimum distance between bounding boxes.
+For two boxes $b_{i}=(x_{1}^{i},y_{1}^{i},x_{2}^{i},y_{2}^{i})$ and $b_{j}=(x_{1}^{j},y_{1}^{j},x_{2}^{j},y_{2}^{j})$, the minimum distance is computed from the horizontal and vertical distances $d_{h}$ and $d_{v}$:
+$d_{\text{min}}(b_{i},b_{j})=\sqrt{d_{h}^{2}+d_{v}^{2}}$
+where $d_{h}=\min(|x_{2}^{i}-x_{1}^{j}|,|x_{2}^{j}-x_{1}^{i}|)$ if boxes don’t overlap horizontally, and $d_{h}=0$ otherwise.
+Similarly, $d_{v}=\min(|y_{2}^{i}-y_{1}^{j}|,|y_{2}^{j}-y_{1}^{i}|)$ if boxes don’t overlap vertically, and $d_{v}=0$ otherwise.
+Two boxes are considered adjacent if $d_{\text{min}}(b_{i},b_{j})\leq\tau$ where $\tau$ is a threshold we choose to be 15 pixels.
+The results are fed into the following graph-based component detection.
 Adjacency Graph Construction.
@@
 This approach effectively reduces element fragmentation while preserving spatial relationships, enabling more accurate element-level descriptions for our hierarchical framework.
-Algorithm 1 provides the complete implementation details.
+Algorithm 1 (at the end of the appendix) provides the complete implementation details.
 B.2 Answer Matching
-To determine answer equivalence, i.e. if two agents agree upon their opinions, we do fuzzy string matching between each pair of answers $a_{1},a_{2}$ using Normalized Levenshtein Similarity (NLS):
+To determine answer equivalence, i.e., if two agents agree upon their opinions, we do fuzzy string matching between each pair of answers $a_{1},a_{2}$ using Normalized Levenshtein Similarity (NLS):
 $\text{NLS}(a_{1},a_{2})=\left(1-\frac{\text{editdistance}(a_{1},a_{2})}{\max(|a_{1}|,|a_{2}|)}\right),$ (5)
-where $\text{editdistance}(a_{1},a_{2})$ computes the Levenshtein distance (Lcvenshtcin, 1966) between the answers after tokenization.
+where $\text{editdistance}(a_{1},a_{2})$ computes the Levenshtein distance (Levenshtein et al., 1966) between the answers after tokenization.
 Two answers are considered equivalent if $\text{NLS}(a_{1},a_{2})\geq 0.75$.
-Appendix C Experimental Setup
+Appendix C Extended Experimental Setup
 C.1 Dataset Descriptions
@@
 Each slide deck corresponds to multiple questions.
-The license is availble at  [^1] https://github.com/nttmdlab-nlp/SlideVQA/blob/main/LICENSE.
-  • InfoVQA (Mathew et al., 2022) is a dataset that feature questions that require reasoning and arithmetic skills.
+The license is available at  [^1] https://github.com/nttmdlab-nlp/SlideVQA/blob/main/LICENSE.
+  • InfoVQA (Mathew et al., 2022) is a dataset that features questions that require reasoning and arithmetic skills.
 It contains over 30,000 questions with over 5,400 images.
@@
 C.3 Implementation Details
-We used EasyOCR [^2] https://github.com/JaidedAI/EasyOCR and Docling (Team, 2024) for detecting textual and visual elements, respectively.
+We used EasyOCR [^2] https://github.com/JaidedAI/EasyOCR and Docling (Auer et al., 2024) for detecting textual and visual elements, respectively.
 Whenever applicable, the answers are generated using a temperature of 0.0 to ensure deterministic results.
+Appendix D Additional Experiments
+D.1 InfoVQA with Open-Source Backbones
+Table 5 reports InfoVQA performance when SlideAgent is paired with open-source MLLMs, complementing the proprietary-model results in the main paper.
+Model Overall Num F1
+Raw LLMs
+Llama-3.2-11B-Vision-Instruct 55.4 49.6 67.9
+Phi-3-vision-128k-instruct 53.4 43.1 88.7
+Qwen2.5-7B-Instruct 80.1 73.0 96.4
+Qwen2.5-32B-Instruct 72.1 62.0 94.8
+llava-1.5-7b-hf 25.1 10.6 83.9
+llava-1.5-13b-hf 20.2 8.0 84.2
+llava-v1.6-mistral-7b-hf 32.6 22.6 79.9
+llava-v1.6-vicuna-13b-hf 34.9 24.8 80.9
+InternVL3-8B 66.7 57.7 85.2
+Multimodal RAG and Agentic Methods
+ViDoRAG 67.1 59.2 85.9
+SlideAgent 75.4 66.5 92.1
+Impr. +8.7 +8.8 +6.8
+Table 5: Performance comparison among open-source multimodal models and agentic methods.
+D.2 Generalizability Across Models
+SlideAgent’s framework is model-agnostic and provides consistent improvements when applied to stronger base models.
+Table 6 shows results with recent state-of-the-art models.
+With Gemini-2.5-Flash, SlideAgent achieves 90.5% overall accuracy (+6.7 points), 87.6% numeric (+9.3 points), and 93.6% F1 (+1.8 points) on SlideVQA, outperforming all models in Tables 2-4.
+With Qwen2.5-VL-7B (released Feb 2025), SlideAgent achieves 80.4% overall (+0.9 points), 73.8% numeric (+3.3 points), and 94.5% F1 (+0.2 points).
+Model Overall Num F1
+Gemini-2.5-Flash on SlideVQA
+Raw Model 83.8 78.3 91.8
++ SlideAgent 90.5 87.6 93.6
+Improvement +6.7 +9.3 +1.8
+Qwen2.5-VL-7B-Instruct on SlideVQA
+Raw Model 79.5 70.5 94.3
++ SlideAgent 80.4 73.8 94.5
+Improvement +0.9 +3.3 +0.2
+Table 6: SlideAgent with recent state-of-the-art models. Gemini-2.5-Flash + SlideAgent achieves best overall performance (90.5%), establishing state-of-the-art on SlideVQA.
+D.3 Generalization to Other Document Types
+To validate that SlideAgent’s hierarchical approach generalizes beyond slide decks, we evaluated on DocVQA (Mathew et al., 2021), a widely-used benchmark for document understanding with scanned documents, forms, and reports (Table 7).
+With GPT-4o as base model, SlideAgent achieves 94.5% overall accuracy (+0.9 points), 91.6% numeric (+2.5 points), and 97.5% F1 (+0.3 points) compared to the baseline.
+With InternVL3-8B, SlideAgent achieves 94.7% overall (+0.5 points), 93.4% numeric (+0.9 points), and 96.5% F1 (+0.5 points).
+Model Overall Num F1
+GPT-4o as base model
+GPT-4o 93.6 89.1 97.2
+SlideAgent 94.5 91.6 97.5
+Improvement +0.9 +2.5 +0.3
+InternVL3-8B as base model
+InternVL3-8B 94.2 92.5 96.0
+SlideAgent 94.7 93.4 96.5
+Improvement +0.5 +0.9 +0.5
+Table 7: Performance on DocVQA, demonstrating generalization beyond slide decks to scanned documents, forms, and reports.
+D.4 Computational Cost and Scalability
+An important aspect of SlideAgent’s design is that computational cost largely depends on the retriever used, which can be swapped based on latency requirements.
+Table 8 compares different retriever configurations versus single-shot GPT-4o.
+Knowledge Construction (One-time, Amortized)
+OCR/layout parsing takes $2.54\pm 0.32$ seconds per page (parallelized).
+Generating hierarchical knowledge ($\mathcal{K}_{g},\mathcal{K}_{p},\mathcal{K}_{e}$) takes 51.5 seconds total for a typical slide deck, using $1544\pm 36$ MB memory.
+This is a one-time cost per document; the generated knowledge acts as a “cache” reusable across multiple queries.
+Retrieval & QA (Per-Query)
+[…]
+Appendix E Use of AI Assistants
+We used AI assistants to improve the paper’s presentation, including language polishing, clarity, and organization.
+AI assistants also provided support for data cleaning and processing.
+All technical experiments, analyses, and final content were verified by humans.
+Model Latency (s) Tokens/sec Memory (MB) Input Tokens
+Raw Model 9.12 $\pm$ 1.12 0.76 $\pm$ 0.27 1253 $\pm$ 35 15420 $\pm$ 2
+SlideAgent (BM25) 9.44 $\pm$ 4.51 0.73 $\pm$ 0.61 1317 $\pm$ 32 1330 $\pm$ 212
+SlideAgent (SFR) 19.17 $\pm$ 13.74 0.13 $\pm$ 0.05 2495 $\pm$ 7 1554 $\pm$ 342
+Table 8: Computational cost comparison on SlideVQA.
+SlideAgent achieves >90% token reduction (15420 $\rightarrow$ 1330 tokens) while maintaining competitive latency.
+The one-time knowledge construction (51.5s total, 1544 MB) is amortized across multiple queries.
+Appendix F Ethical Considerations
+Data Privacy, Consent, and Intellectual Property.
+Visual documents such as those processed by SlideAgent may contain sensitive business or personal data.
+Ensuring compliance with privacy regulations (e.g., GDPR, CCPA) and obtaining appropriate consent are essential.
+Organizations should also respect intellectual property rights and establish policies that balance knowledge sharing with fair use.
+Content Reliability and User Responsibility.
+Like other LLM-based systems, SlideAgent may inherit biases or produce inaccurate content.
+While it enhances question answering, outputs should be verified by human judgment, particularly in sensitive or high-stakes scenarios.
+Clear user guidance and validation practices can help ensure responsible use.
+Figure 9: Performance comparison among variants of SlideAgent with base model InternVL3-8B.
 Family Model Parameters Knowledge Cutoff Release
@@
 Table 9: Overview of models used in the experiments, including their family, model name, parameter size, knowledge cutoff date, and release date.
+Notation Description
+$p_{i},v_{i}$ A slide page and its visual content
+$\mathcal{P}=\{p_{1},p_{2},\dots\}$ Set of pages in the slide deck
+$q$ User query
+$\hat{Q}$ Set of subqueries generated from the original query $q$
+$a$ The final answer generated by SlideAgent
+$e_{j}$ Individual element in a slide, consisting of visual content, bounding box, and type
+$b_{j}$ Bounding box coordinates of an element within a slide
+$t_{j}$ Type of the element (e.g., text, chart, image, table)
+$\mathcal{K}_{g},\mathcal{K}_{p},\mathcal{K}_{e}$ Global, page, and element knowledge
+$\mathcal{M}_{g},\mathcal{M}_{p},\mathcal{M}_{e}$ Global, page, and element agent
+$\mathcal{R}$ Retrieval function that fetches relevant pages or elements based on subqueries
+$\phi(\cdot)$ Answer synthesizer that combines reasoning from all agents to generate the final answer $a$
+[…]
+$h_{g},h_{p},h_{e}$ Answers and reasoning from global, page, and element agents
+Table 10: Mathematical notation used throughout the paper.
+Model SlideVQA TechSlides FinSlides
+Overall Num F1 Overall Num F1 Overall Num F1
+Raw Models
+Gemini 2.0 86.3 81.0 90.3 59.7 60.0 59.6 78.1 77.8 88.9
+Gemini 2.5 89.0 85.7 93.2 61.5 65.0 59.5 76.6 76.4 83.3
+Gemini 2.5-lite 81.8 75.5 90.1 56.3 55.0 57.0 78.1 78.4 66.7
+Claude 4.1 85.7 82.4 89.7 58.0 77.5 48.3 52.6 52.0 60.8
+Claude 3.5 58.2 64.0 50.9 55.8 83.7 42.5 47.8 48.0 40.3
+GPT-4o 79.4 71.9 86.4 64.5 77.1 58.0 83.0 84.3 80.1
+Multimodal RAG and Agentic Methods
+ViDoRAG 81.8 73.8 87.0 65.8 78.0 58.7 84.2 84.7 81.5
+SlideAgent 87.1 84.4 90.6 68.7 82.5 61.5 85.8 85.9 85.6
+Impr. +7.7 +12.5 +4.2 +4.1 +5.4 +3.5 +2.8 +1.5 +5.5
+Table 11: Performance comparison of baselines and proprietary models on SlideVQA, TechSlides, and FinSlides, assuming access to ground-truth pages containing the correct answer.
+All baseline methods use GPT-4o for question-answering.
+Model SlideVQA TechSlides FinSlides
+Overall Num F1 Overall Num F1 Overall Num F1
+Raw Models
+Llama-3.2-11B-Vision-Instruct 44.6 52.1 34.6 47.1 62.8 39.3 39.1 39.2 33.5
+Phi-3-vision-128k-instruct 78.3 69.1 91.6 53.9 67.4 47.2 63.8 63.7 65.1
+Qwen2.5-VL-7B-Instruct 85.1 77.7 95.3 57.7 69.8 51.8 52.7 52.0 77.8
+Qwen2.5-VL-32B-Instruct 87.4 82.6 93.7 49.6 62.8 43.2 69.5 69.6 65.1
+llava-1.5-7b-hf 42.9 27.9 79.9 24.6 15.0 39.4 14.2 14.4 20.9
+llava-1.5-13b-hf 46.7 29.5 83.1 29.2 17.5 46.6 23.8 20.3 41.2
+llava-v1.6-mistral-7b-hf 59.0 45.8 84.2 36.2 40.0 34.0 16.0 15.2 21.3
+llava-v1.6-vicuna-13b-hf 62.3 48.2 87.1 54.7 62.5 49.5 35.2 30.7 67.6
+InternVL3-8B 73.3 65.4 85.7 58.4 72.1 51.5 56.3 55.9 65.6
+Baseline methods based on InternVL3-8B
+ViDoRAG 76.3 68.1 89.9 61.3 75.1 53.9 58.4 58.7 67.3
+SlideAgent 82.8 75.3 93.3 64.6 79.5 58.0 62.8 62.6 68.1
+Impr. +9.5 +9.8 +7.6 +6.2 +7.4 +6.4 +6.5 +6.7 +2.5
+Table 12: Performance comparison of baselines and proprietary models on SlideVQA, TechSlides, and FinSlides, assuming access to ground-truth pages containing the correct answer.
+All baseline methods use InternVL3-8B for question-answering.
+Input Initial global knowledge $\mathcal{K}_{g}^{(0)}$ and concatenated page-level knowledge $\mathrm{concat}(\mathcal{K}_{p}^{1},\ldots,\mathcal{K}_{p}^{|\mathcal{P}|})$.
+Instruction Rewrite the global knowledge from the full page-level evidence.
+Use $\mathcal{K}_{g}^{(0)}$ only as a weak prior for terminology and early context.
+Incorporate complete document coverage and narrative flow, including conclusions, summaries, and next steps that appear in later pages.
+Update Rule Regenerate each field rather than appending snippets.
+If later-page evidence contradicts or refines an earlier hypothesis, overwrite the earlier field with the better-supported document-level interpretation.
+Output Fields Title, objective, structure overview, key insights, audience, and tone.
+Table 13: Prompt template for the global refinement operator. The operator performs a single-pass fieldwise rewrite from full-document page evidence.
+You are given a complete slide deck consisting of multiple slides. Your task is to synthesize a concise, high-level summary of the overall message, structure, and purpose of the deck.
+### Cumulative Summary of the Preceding Slides {Cumulative Summary}
+### Response format Please respond with a markdown string that follows the following format: Title <<Concise Explicit / Inferred Title of the Presentation>>
+Objective <<What is the presentation trying to achieve? (e.g., inform, persuade, pitch, propose)>>
+Structure Overview Slide 1: <<Brief description of the slide>> Slide 2: <<Brief description of the slide>> Slide N: <<Brief description of the slide>> Key Insights <<Major takeaway 1>> <<Major takeaway 2>> Audience <<Intended audience type (e.g., executives, investors, engineers)>>
+Tone <<Overall tone: e.g., persuasive, analytical, optimistic, urgent>>
+Table 14: Prompt template for the global agent to generate comprehensive slide deck summaries, including title, objective, structure, key insights, audience, and tone.
+Element Type <<text | image | chart | table | icon | button | etc.>>
+Position on Slide <<e.g., top-right, centered, below title>>
+Verbatim Content <<if text, give the literal string; else describe the visual>>
+Semantic Role <<What is the element trying to do or communicate?>>
+Functional Purpose <<Its practical function within the slide (e.g., emphasize point, guide attention, show evidence, support action)>>
+Relation to Slide <<How does it connect to or support the slide’s overall message?>>
+Inferred Importance <<how central is this element to the slide? Answer with low, medium, or high>>
+Table 15: Element-level slide description prompt format used by the element agent to generate detailed annotations for each visual component.
 Algorithm 1 Graph-based depth-first search for merging fragmented elements into coherent semantic units while preserving spatial layout.
@@
 30:return $\tilde{B}$
-Appendix D Additional Experiments
-D.1 Ablation Studies (Cont’d)
-Element-level Reasoning Ensures Precision
-Ablating the element agent causes a moderate yet consistent decline (–4.6 overall for GPT-4o; –6.3 for InternVL3-8B), especially numeric questions.
-Even with perfect retrieval, omitting fine-grained texts and layouts weakens factual grounding.
-Impact of Global Thematic Guidance
-Removing the global agent yields the smallest drop (–2.8 GPT-4o; –3.7 InternVL3-8B), as lower-level agents already embed partial global context via $\mathcal{K}_{p}$ and $\mathcal{K}_{e}$ in knowledge construction (Eq.
-1/2).
-However, responses become less aware of overarching themes.
-Subquery Generation Strengthens Retrieval.
-Removing subqueries causes larger losses under retrieval (–5.0 GPT-4o; –11.3 InternVL3-8B) than with ground-truth pages (–2.9 to –4.9).
-Visual-aware subqueries markedly improve retriever accuracy by aligning textual intent with visual semantics–especially beneficial for weaker open-source encoders.
-Appendix E Future Work
-Element Parsing.
-SlideAgent relies on advanced text and layout detection tools, which may struggle with visually complex or low-contrast designs (e.g., white charts on white backgrounds).
-Enhancing robustness in parsing could further improve overall reliability.
-Modeling Element Relations.
-SlideAgent currently treats elements independently.
-Extending it to explicitly model inter-element relations–for instance, via graph structures linking textual and visual components–offers a promising direction for capturing richer semantics, though at higher computational cost.
-User-Centric Interaction.
-While SlideAgent primarily enhances reasoning and retrieval for slide understanding, future work could make such agentic frameworks more user-centric through seamless integration into real-world user behaviors Wang et al.
-(2025b).
-Lightweight feedback mechanisms (e.g., contextual pop-ups or in-situ clarifications) could further improve interpretability and interaction without disrupting natural workflows or requiring excessive user input.
-You are given a complete slide deck consisting of multiple slides. Your task is to synthesize a concise, high-level summary of the overall message, structure, and purpose of the deck.
-### Cumulative Summary of the Preceding Slides {Cumulative Summary}
-### Response format Please respond with a markdown string that follows the following format: Title <<Concise Explicit / Inferred Title of the Presentation>>
-Objective <<What is the presentation trying to achieve? (e.g., inform, persuade, pitch, propose)>>
-Structure Overview Slide 1: <<Brief description of the slide>> Slide 2: <<Brief description of the slide>> Slide N: <<Brief description of the slide>> Key Insights <<Major takeaway 1>> <<Major takeaway 2>> Audience <<Intended audience type (e.g., executives, investors, engineers)>>
-Tone <<Overall tone: e.g., persuasive, analytical, optimistic, urgent>>
-Table 9: Prompt template for the global agent to generate comprehensive slide deck summaries, including title, objective, structure, key insights, audience, and tone.
-Element Type <<text | image | chart | table | icon | button | etc.>>
-Position on Slide <<e.g., top-right, centered, below title>>
-Verbatim Content <<if text, give the literal string; else describe the visual>>
-Semantic Role <<What is the element trying to do or communicate?>>
-Functional Purpose <<Its practical function within the slide (e.g., emphasize point, guide attention, show evidence, support action)>>
-Relation to Slide <<How does it connect to or support the slide’s overall message?>>
-Inferred Importance <<how central is this element to the slide? Answer with low, medium, or high>>
-Table 10: Element-level slide description prompt format used by the element agent to generate detailed annotations for each visual component.
-Model SlideVQA TechSlides FinSlides
-Overall Num F1 Overall Num F1 Overall Num F1
-Raw Models
-Gemini 2.0 86.3 81.0 90.3 59.7 60.0 59.6 78.1 77.8 88.9
-Gemini 2.5 89.0 85.7 93.2 61.5 65.0 59.5 76.6 76.4 83.3
-Gemini 2.5-lite 81.8 75.5 90.1 56.3 55.0 57.0 78.1 78.4 66.7
-Claude 4.1 85.7 82.4 89.7 58.0 77.5 48.3 52.6 52.0 60.8
-Claude 3.5 58.2 64.0 50.9 55.8 83.7 42.5 47.8 48.0 40.3
-GPT-4o 79.4 71.9 86.4 64.5 77.1 58.0 83.0 84.3 80.1
-Multimodal RAG and Agentic Methods
-ViDoRAG 81.8 73.8 87.0 65.8 78.0 58.7 84.2 84.7 81.5
-SlideAgent 87.1 84.4 90.6 68.7 82.5 61.5 85.8 85.9 85.6
-Impr. +7.7 +12.5 +4.2 +4.1 +5.4 +3.5 +2.8 +1.5 +5.5
-Table 11: Performance comparison of baselines and proprietary models on SlideVQA, TechSlides, and FinSlides, assuming access to ground-truth pages containing the correct answer.
-All baseline methods use GPT-4o for question-answering.
-Model SlideVQA TechSlides FinSlides
-Overall Num F1 Overall Num F1 Overall Num F1
-Raw Models
-Llama 3.2 11B 44.6 52.1 34.6 47.1 62.8 39.3 39.1 39.2 33.5
-Phi3 78.3 69.1 91.6 53.9 67.4 47.2 63.8 63.7 65.1
-Qwen2.5 7B 85.1 77.7 95.3 57.7 69.8 51.8 52.7 52.0 77.8
-Qwen2.5 32B 87.4 82.6 93.7 49.6 62.8 43.2 69.5 69.6 65.1
-LLaVA 1.5 7B 42.9 27.9 79.9 24.6 15.0 39.4 14.2 14.4 20.9
-LLaVA 1.5 13B 46.7 29.5 83.1 29.2 17.5 46.6 23.8 20.3 41.2
-LLaVA 1.6 7B 59.0 45.8 84.2 36.2 40.0 34.0 16.0 15.2 21.3
-LLaVA 1.6 13B 62.3 48.2 87.1 54.7 62.5 49.5 35.2 30.7 67.6
-InternVL3 8B 73.3 65.4 85.7 58.4 72.1 51.5 56.3 55.9 65.6
-Baseline methods based on InternVL3-8B
-ViDoRAG 76.3 68.1 89.9 61.3 75.1 53.9 58.4 58.7 67.3
-SlideAgent 82.8 75.3 93.3 64.6 79.5 58.0 62.8 62.6 68.1
-Impr. +9.5 +9.8 +7.6 +6.2 +7.4 +6.4 +6.5 +6.7 +2.5
-Table 12: Performance comparison of baselines and proprietary models on SlideVQA, TechSlides, and FinSlides, assuming access to ground-truth pages containing the correct answer.
-All baseline methods use InternVL3-8B for question-answering.
+Figure 10: Sample global knowledge $\mathcal{K}_{g}$ generated by the global agent, showing document-level summary, objectives, and narrative structure.
+Figure 11: Sample page-level knowledge representation $\mathcal{K}_{p}^{i}$ generated by the page agent, capturing slide-specific content and cross-slide relationships.
+Figure 12: Sample element-level knowledge representation $\mathcal{K}_{e}^{j}$ generated by the element agent.